A Complete Bibliography of Publications in Scientometrics: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

20 [Gup89]. 80 [Gup89]. ^{(}R) [SS85]. R(r) = a.r − b [Sen89b].

-Rule [Gup89].

/ [Gup89].

1870AD [She80]. 18th [She82]. 1970s [SG86]. 1984 [Ano85b, Mor85a]. 1985 [Ano87a]. 1986 [Ano87b]. 1987 [Ano88b, Ano89b, Gra89a, Van88]. 1988 [Ano89a, Van89a]. 19th [She82].

80/20 [Egg86].

A. [BIM^{+}87]. A.D. [Sim80]. academic [CGH^{+}88, Kea91, Le 80, LM85, Yan82, vHvVLD83]. accept [BGS89a]. acceptance [Gor80]. accounts [Hur81]. achievements [Her88, Kea91]. acidification [LBCW88]. Acquired [SFL89]. acrobats [Nal87]. across [Hop84]. activities [FD86, Hai83, Hai86a, Hai86b, Pav85, Pav88]. activity [BJ81, BD87, Fra80, MGFA87, PABK88, Sim80, Vin88c, Yut86]. adaptation [Kre87]. adaptive [De 87]. additivity [Hai86a, Hai86b]. Address [Mor85a]. adequate [AM89]. adjusting [Lin82, LM82b, LM82a]. advantage [Law87]. Advisory [Ano82a]. Africa [Dav83, Pou89]. against [Hai88, Kun88a]. age [Han89, Ove88, VD87a, VD87b]. ages [HG85]. aging [Ave85, Chr85, Gie81, MT89, Mot89]. Agricultural [Bal85a, Bal85b, Bal85c, Mor87a, Pou89]. aid [CF82]. AIDS [SFL89, Coh84]. alcohol [SK86]. alive [Fra84a]. Allocation [JKW87]. along [KPV86]. Alphonse [Sza85]. America [KPV86]. American [FP88, FKM81, Oro81, Oro85, Ric84]. among [Alo85, Coz85, FP88, Har86, Moe89, Ric84]. analysed [VP89]. Analysis [NG86, SK89, Yan82, Ada81, Ave80, BCC89, BJ81, CC82, CG80, Cou89, CIM86, Dav83, De 87, DT85, FHNR81, Gar85, GR88, GG88, HS89, Hai82a, Hai86a, Hai86b, Hje82, Lan85, LBCW88, LKO80, Lev86, Mal89, MC85, MT89, Mid83, MBFV85b, Mor86b, Okr87, Oro81, Raj83, SZB83, SZTB84, Sha89, Sha87, SHC89, SMKG86, Sim80, Sim86b, SG80, SBN84, SKWK80, TDV87, Tro87, Yan81, Zuc87, de 81a, de 81b]. analytical [DHQ89]. analyzing [Nom82a]. Announcement [Ano84a]. answers [Mor87b]. Applicability [EC87]. Application [Gup89, DS89, Eto88, Kre83b, Kry88, MBFV85a, Sen86a]. applications [SG84]. Applied [Mor85b, Nor87, Alo85, Fra84a, Mor89a, Mor89b]. approach [CL87, Lib85, POV86, Sen89a, Tod89]. Approximating [RRJ83]. approximation [CC82]. Arab [Mor80]. archive [Coz85]. area [BCC89]. areas [JNR86]. Article [Sha87, Hai88, Mar81b, Sni84]. Article-by-article [Sha87]. articles [BCC89, Dia85, MT89, Nom82b, Oro85, SB86b]. artifacts [MM87]. Asian [DE89]. ask [Mor87b]. aspects [LvE81, PL86]. aspirations [Chu85]. ASR [SDH82]. Assessing [BGS89a, Lin88, NZDD89, Lea85, SGB87, Luu89]. assessment [Ami85, Mor85b, Pou89, PABK88, SB86a, Sim84, SH83, Vin86]. assessments [BGS89a]. Astronomy [Her86, AH85, Gie81, Her88]. astrophysics [AH85]. attempt [Vin88a]. attendance [Lyo80, SZB83]. attitudes [LC80]. Aurel [Bal87]. Author [SB86b, McC83, RRJ83, Swa85, Vel87]. authored [Dia85, SB86b]. authors [Her88, PP85]. Authorship [Bal85b, SR89, Ste82, Gor80, Hef81, Lin82, LM82b, LM82a, POV86]. avian [Kin88]. avoided [Gil89]. Avramescu [Bal87]. Award [Ano85b, Ano87a, Ano87b, Ano88b, Ano89a, Ano89b, Fra87, Gra89a, Mor85a, Mor87c, Van88, Van89a, Ano85a].

B. [Kea91]. backs [Lin87b]. Bank [Mor86b]. Barnaby [BZ85]. based [AM89, Bon85, Pin80, SG83, Sla83, TA86]. bases [Ste87]. basic [Bow80, Gua86, Kid88, Nor87, Rie80, SR89, TK81]. basis [BdSP81]. BCS [Nad81]. be [MBFV85a]. becoming [LvE81, NN85]. before [Mor87b]. behavioral [Rie80]. behaviors [Ono88]. behaviour [LD86]. behavioural [Lan85, NZDD89]. beholder [Vel87, Mor87a]. Belgium [FD86]. beliefs [FKM81]. believe [Nar87]. best [Gil89]. bestowers [Sni84]. better [Lin88]. Between [CA89, Gor80, Hai89, Hal87, KAA^{+}88, Lan86, Le 80, Ley89a, Sen89b, She80, VJL89, VD87a, VD87b]. Beyond [Chu85]. Bias [Shr85]. Bibliographic [Sen89a, Sha89, SMKG86, Sma87, Ste87]. bibliographies [Ano81a, Ano81b, Ano81c, Ano81d, Ano81e]. Bibliography [Hje82, DH89, Sch83a, Sch83b, Sch83c, Sch84a, Sch84b, Sch84c, Sch85a, Sch85b, Sch86, Sch87, Sch88a, Sch88b, Sch89a, Sch89b]. Bibliometric [CGH^{+}88, CIM86, GR88, Moe89, Vin88b, Yut86, AG85, Kin88, Law86, LKO80, Lev86, MBFV85a, MBFV85b, NV87b, NZDD89, NV89, Nor87, Ono88, PCJ88, RL85, Sen86a, Tod89, Vin88a, Woo88]. Bibliometrics [PT88, SK86, Bro87, Hje82, Kry88, Per86, Sit87]. bibliometry [Kea91]. big [AG85, CM85, Gar85, IM85, PK87]. biochemical [Gup89, Mar81b]. biochemistry [Sen86a]. biologists [She82]. biology [BN82, Nor87]. biomedical [Chr85, Coh80, Coh81, MC85]. biophysical [Sen85]. biotechnologists [Ned88]. biotechnology [RC84]. Board [Ano82a]. Bonitz [OV89b]. Book [Bro88, BN84, Bru81, EM83a, GMS84, Gri81, HFG82, Inh81, IS82, Jac87, KS89a, KDYR87, LM88, McA81, ML80, MM81, MPMG82, Mor83, MM85, Mor85c, Mor88, NM86, RFH^{+}83, RMD86, SM83, SMR84, SZTB84, YT89]. Bradford [Bon80, CL87, Eto84, EC87, Eto88, KLC84, Mor81, Nar89, PL86, Sen89b]. Brannigan [Sim86b]. Braun [Ano87b, Mor87c]. Brazilian [Mor87a]. Britain [Kea91]. British [BGS89a, CGH^{+}88, Pav88, RM81].

calculation [BdSP81, Egg87]. Call [Ano81a, Ano81b, Ano81c, Ano81d, Ano81e]. can [Vel87]. Canada [CF82, Bin81]. Canadian [RM81]. cancer [Luu89]. Candolle [Sza85]. career [LM85]. Carlo [Sim86a]. Case [NG86, Bin81, FKM81, GR88, HS89, Hai88, Hal87, Kun88a, LPP^{+}86, Per83a, Sni86, Swa88, TK81, VJL89, Van89b, PABK88]. Casting [Sni84]. causal [Hop84, Sen86b]. causes [ER88, She80]. center [HG85]. centers [CA89]. centuries [She82]. century [Hal81, Her86, Her88]. CERN [IM85]. chance [Sim86b]. change [LBCW88, SKWK80]. Changes [Mar81b, Ned88]. characteristic [SGB87]. characteristics [Mor81]. characterizing [Vin88c]. charts [BGS89c, SB86a]. Chemistry [Ste82, BN82, BGS87b, BGS88b, DNW80, Gua86, NV89, PP85, SZTB84, ZSB82a]. Chinese [FN87]. Choosing [Dan87]. Citation [Ave85, De 87, Gre83, Gre84, Hai82a, Mar81a, MT89, Mid83, Nad81, NV87a, Pin80, SG80, SS85, Vla85, Yan81, ZSB82a, Zuc87, AM89, BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89b, BGS89c, BGS89d, CC82, CG80, DNW80, FHNR81, Gar85, Hje82, Lan85, Ley87a, Ley89b, Lin89, Luu89, MG89, MvR86, Ned85, NR88, Per83b, Plo89, PABK88, RBBH89, RRJ83, RMM87, SG83, SB86a, SGB87, Sha87, Sla83, SSG85, SB86b, SKWK80, VJL89, Vel86, Vin87]. citation-based [AM89]. Citations [AL89, PCJ88, Bal85c, BdSP81, Coz89, Dia85, Kre83b, Kre83a, Rou87, RM81, SS85, SSG85]. cited [Ave85, LPP^{+}86, Nom82b, Per83a, PL87]. citing [Nom82b, PL87, PP85]. class [SG84]. classics [MG89]. classification [Mor86a, Per83b]. classifications [SH83]. classifying [Vin88a]. climate [SK89]. clinical [SR89]. Cluster [Sha89, SG80]. Clustering [SS85, SSG85, Sla83]. Co [RC84, GG88, Kor89b, LBCW88, Nad81, SG80, SS85, SSG85, SKWK80, TV89, VP89]. co-citation [Nad81, SG80, SKWK80]. co-citations [SS85, SSG85]. co-modelling [Kor89b]. co-nomination [GG88]. co-subfield [VP89]. Co-word [RC84, LBCW88, TV89]. Cocitation [Hop84, McC83, MVVL85]. coefficient [ABT88]. Cognitive [Oro81, Koc83, RC84]. collaboration [ABT88, Gor80, Hef81, POV86, SB86b, Van81]. collaboration/scientific [POV86]. Collaborative [ABT88]. Collagen [SG86]. collection [BGS89c]. colleges [Chu85, Mul85]. comment [Kea91]. Comments [Col80, Fra87, Gra89a, Man80a, Mor87c, Ros80, Van88, Van89a, de 80, SDH82, Sni84]. commitments [Vel87]. common [Nom82b]. Communication [FD86, Lib85, Man80b, Sha81, Swa88, TA86, VR88]. communities [DE89]. comparable [SW83]. Comparative [DT85, NK88, BN82, CCM88, Kea91, MBFV85b, NZDD89, NV89, PNN87, SB86a, vRH87]. compare [AM88, McC80]. Comparing [PCJ88, PL87]. Comparison [Mid83, KAA^{+}88, MG89, NV87b, Ono88, TV89, Woo88]. comparisons [IM89, SG83, ST86]. COMPENDEX [MGFA87]. competence [DT85, NG89]. competitive [JKW87]. complex [Alo85, Kre83b, Zuc87]. composition [CN80]. comprehensive [SGB89]. computer [BdSP81]. computerized [Hal89, JKW87]. Concentration [CF82, EM83b, Eto84, Mal81]. conceptualization [HMM88]. Condensed [Tod83]. conference [Lyo80]. Conferences [NG86, KL89]. conjectures [Hai88, Kun88a]. consensus [KCB82, Oro81]. consequences [Oro87, Vin88b]. considered [MBFV85a]. considering [Ono88]. constant [HZ84]. consumer [Kea91]. Consumption [SE82]. contemporary [Sim84]. context [MT89, SG80, Vel86]. Contextual [SHC89]. contribution [Vin88c]. control [Dia87]. Cooperation [Kre85, Kre87]. cooperative [Bin81]. Corollaries [Hai86b]. Correction [Ano84b, Bli82]. correlates [Law86, SDH82]. correlation [Sen89b]. correspondence [TDV87]. Council [CF82, Mcg88]. count [Coz89]. countries [BGS87c, BGS88c, BS88, CCM88, Chr85, Fra80, Lan85, Lan86, Mor85b, NK88, Poi84, PABK88, PNN87, SZB85, SGB89, Vin88c]. country [AG85, Vel86]. Counts [FP81, Lin89, Ned85, RRJ83]. courses [Ano82b]. coverage [Kid89]. Creativity [Pur88, Pur89, Han89, RMP83]. critical [Gor80]. Croatia [PABK88]. Croatia/Yugoslavia [PABK88]. Cross [KM80, PC85, ST86]. cross-disciplinary [PC85]. Cross-national [KM80, ST86]. crystallography [CIM86]. Cuba [LPP^{+}86, Mor89a]. cultural [Pur89]. cultures [Roc83]. cum [NV87b, NV89]. cumulative [Law87]. current [Sch83a, Sch83b, Sch83c, Sch84a, Sch84b, Sch84c, Sch85a, Sch85b, Sch86, Sch87, Sch88a, Sch88b, Sch89a, Sch89b]. currents [CIM86]. curriculum [Sit87]. cybernetic [She80]. cycle [Dia84, Dia87].

data [BGS89d, Hai86a, Hai86b, Has81, RBBH89, Rou89, Ste87, Vla85]. database [BGS89d]. databases [Hal89, MG86]. datafiles [SGB89]. day [CM81]. dealing [Bal85b]. death [Gre83, Gre84]. decade [Hop84]. December [Mor85a]. decision [DHQ89]. decline [BGS89a, Ove88]. definition [Bro87]. deflection [ER88]. degree [ABT88, Ami85, JSCL82]. Department [Por88]. Departments [Cre88, Gil89]. dependent [MBFV85a]. Derek [Ano85b, Ano87a, Ano87b, Ano88b, Ano89a, Ano89b, Gra89a, Van88, Van89a, Ano84c, Ano85a, Coz85, Fra87, Gar85, Gri84, Mor85a, Mor87c]. derived [Ono88]. Determinants [Koe82]. Determination [Tod84]. determine [Rou87]. Determining [Gil89]. develop [CGH^{+}88]. developed [Chr85, Fra80, PABK88]. developing [Chr85, Mor85b]. development [Ano84d, Bia87, Ley86, LD87, Mor89a, Mor86b, Por88, Pur88, Raj83, SK89, Tro87]. developments [Van89b]. diagnosis [Mor89a]. dice [Lin88]. Did [KL85]. Differences [Okr87, Coz85, Hur81, Kyv89, Law87, Moe89]. different [Bal85b, ER88]. dilemmas [Nas88]. Dimensions [NK88]. directory [Ano82b]. disciplinary [PC85]. disciplines [FD86, Har86, Hef81, Kot87, Lan85, Lan86, Le 80, vHvVLD83]. discovery [Hal87, KL85, Sim86a]. display [vRH87]. Displaying [MVVL85]. dissertations [Woo88]. distortion [Hai83]. Distribution [KCB82, Tod85, Bal85c, BS82, BGS89d, BdSP81, CdA89, EM83b, EC87, Eto88, GS85, Hai89, Mar81b, Mor81, OV89a, Ono88, PL86, Yab80]. distributions [Hai82b, Hai82c, Hai82d, Hur84, Mok89, SG84, Sim86b, Yab85]. DNA [SG80]. do [AM88, Coz89, Pav88]. Dobrov [Kor89a]. doctorates [NV87b, NV89]. document [MVVL85]. Does [Ove88, McC80]. domination [FP88]. dragons [Kun88b]. droop [ER88]. Dual [POV86]. during [Hal81, Tro87]. Dutch [HS89, Sni86, Van81]. dynamic [SG84]. Dynamics [LS87, VP89, Sma88].

early [Sza85]. earth [Hal81]. easy [NR88]. ecology [HS89]. Econometric [MC85]. economic [Ano84d, BS88, Dia84, LR84, Raj83]. economists [FP88]. economy [Fre80, Ros80]. editor [KM88, Lin87a]. Editorial [Ano82a, Ano83, Bra89, ZSB82b, ZSB82a]. Editors [BR85]. education [TK81]. Effect [NG89, BZ85, Hai83, NR88, OV89b]. effectiveness [RMP83]. Effects [Lan85, LM85]. efficiency [Pok80]. eight [MG86]. electricity [Kry88, Kry89]. electron [VJL89]. elemental [KL85]. Elements [Mor89a]. elite [Shr85]. elites [Mul85]. Emergence [Gua86]. eminence [Sim84]. eminent [FP88]. empirical [Lan85, Lan86, Ono88, Oro81, Per83a, SK86, SH83]. employment [Raj83]. Energy [LKO80]. engineering [BGS88a, KL89, MGFA87]. engineers [Alo85]. English [SK86, Swa85, Yam89]. English-language [Yam89]. entomological [Gup87]. Environment [Mcg88, JKW87]. environmental [LBCW88]. epidemic [SFL89]. equation [SK86, Sen89b]. essence [Nas88]. estimate [NR88]. estimation [CM81, DT85, Nar89, PK87]. ethnomethodology [Hop84]. Eugene [Ano85b, Mor85a]. European [Poi84]. Evaluating [IM85, Mor89b, Ned85, SZB85]. Evaluation [Vin86, BN82, CG80, CIM86, GG88, Kin88, Ned88]. every [Per86]. Evidence [Bon80, KM80, Lin82, LM82a, Mar85]. evolution [Kot87]. exact [Egg87, GS85]. examination [Hur84, Sni86]. example [RC84]. expending [Eto84]. expenditure [Raj89]. experience [Por88]. experiences [Mcg88]. experiment [CGH^{+}88, Gra89b]. experiments [Sim86a]. expert [Koe82]. expertise [DS89]. experts [KCB82]. exploitation [Gie81]. explorations [Swa85]. extreme [Mok89]. eye [Mor87a].

facility [Sen86b]. factor [SE82]. Factors [LPP^{+}86, Bon85, Gro89, MBFV85a]. Facts [MM87, BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89a, BGS89c]. features [Vin88b]. field [BdSP81, DD82, KL89, LKO80, MBFV85a, RMM87, SGB87, Sen86a, Sen89d, Sit87, Sza85, Tod89, VP89, Van89b]. field- [MBFV85a]. fields [Kyv89, MG89, Per83b, SGB89, Shr85]. figures [BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89a, BGS89c]. film [DD82]. finance [Mor86b]. firms [DL85, Eto84]. first [Coz89, Mor85a, RRJ83, Sni84, Swa85]. fisheries [BJ83, BJ84]. flow [Bal86, DD82]. flows [LD86, LD87]. Forecasting [Raj89]. foreign [Hai82a, SW83]. form [Mok89]. Formal [TA86]. formula [Sen89d]. formulation [GS85]. four [DE89, Hef81, Hur81, Lyo80]. frames [Ley86]. framework [Lea85]. Francis [Ano89a, Van89a]. frequency [Hub81, Ono88, Plo89]. fresh [HS89]. fresh-water [HS89]. fuel [SK86]. fuels [Hal87]. function [Coh80, Coh81, Ono88, Qur84]. fundamental [Yab80]. Funded [Hef81, Kea91]. funding [Poi84]. funds [JKW87]. Further [Lin82, LM82a]. future [AH85, IM85]. Fuzzy [DS89, Kun88b].

games [Lyo84]. gap [VJL89]. Garfield [Ano85b, Mor85a]. gatekeepers [ZSB82a]. gatekeeping [ZSB82b]. Gaussian [Hai83, Yab85]. Gedanken [Sim86a]. generalised [Sen89a]. generalized [TA86]. generation [Flo88]. genetics [Bal86, MT89]. Gennady [Kor89a]. geographic [CdA89]. geographical [Hai82a, Mal81]. geoscience [Hal87, Hal89]. giants [Lin87b]. giving [Mor87b]. good [Kea91]. Government [Kea91, Ned88]. Government-funded [Kea91]. Gozinto [Rou87]. grants [CF82]. graphical [vRH87]. greatest [Her88]. Groos [ER88]. Gross [OV89b]. group [Kre85, MvR86, NG89, SH83]. groups [Kre83b, Kre85, Kre87, PNN87, Van81]. Growth [BZ85, JNR86, KS89b, FN87, Hal89, KB81, SFL89, Sen85, Sen89c, Woo88]. guide [KL85].

habits [Lan85]. Haitun [Kun88a]. health [CdA89]. Heat [NG86]. Henry [Ano88b, Van88]. Hierarchical [Sla83]. high [DL85]. highly [Ave85]. historical [Hur81, Sim84]. histories [Vla85]. history [Hur81, Kry88, Kry89]. human [BM86, Ono88]. humanities [NZDD89, Woo88]. Hungarian [BN82, SB81]. Hungary [JHBH88]. hypotheses [MvR86]. hypothesis [Mar85, Mea87, CC87, Law87, MM87, NV87a, Oro85, Sni86, Sni87].

iceberg [SZTB84]. identical [Man80b]. II [Hai86b, Kid89, SSG85, VD87b]. immediacy [Mar85]. immigrants [LR84]. immunodeficiency [SFL89]. Impact [OV89a, Ano84d, Bon85, BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89b, BGS89c, BJ81, Fre80, Nad81, Ned88, Ove88, Plo89, Pok80, Ros80, RM81, SG83, SB86a, SB86b, TK81, Vin88c, vRH87]. implications [SE82, de 81a]. Important [MBFV85a]. improved [Nom82a]. increase [Han89, OV89a]. independence [JSCL82]. Index [SS85, BGS89d, CL87, DNW80, Ley89b, PABK88, SSG85]. indexing [Hje82]. India [Gua86, Raj89]. Indian [GR88, Raj83]. indicator [Plo89, PC85, PABK88, ST86, SW83, ZSB82b]. Indicators [AL89, Ano84d, Bow80, Fre80, Kuh80, LC80, PNN87, ZM80a, Ave80, BS88, Bro80, CGH^{+}88, Dan87, EC87, Gro89, Man80a, McC80, MBFV85a, Nad81, NV87b, NZDD89, NV89, Pav85, Por88, RL85, SZB85, SB86a, SGB89, Vin88a, Vin88c, Yut86, Col80, Ros80]. Individual [Ste82, Law87]. industrial [Gua88, Ned88]. industry [Bia87, EM83b, Gua88, Hal87, SHC89, Por88]. inequality [Okr87]. infant [Gre83, Gre84]. inference [Coh84]. inferred [Gor80]. influence [Sni87, Tod84, Vel87, Zuc87]. Influences [Gro89, Rou87]. influencing [LPP^{+}86]. influential [Oro85]. Infographic [DHQ89]. informatika [Ste85]. Information [Bal86, Kor89b, Sha81, Alo85, BS82, DD82, Flo88, Kun87, LD86, Man80b, Mar81a, Mor89a, Sit87, Ste85]. innovation [EM83b, Gie81]. innovative [Pav85]. input [Dan87, SBN84]. input-output [SBN84]. insecticide [RL85]. INSPEC [MGFA87]. Inspirations [Chu85]. Instead [Ano85c]. Institute [Bin81]. institutes [SB81]. institution [Coh80, Mor86b, SMKG86]. Institutional [Dav83, Kin88]. institutions [CGH^{+}88, Coh81, Kor89b, Yan82]. instrumental [VJL89]. intellectual [Zuc87]. intelligence [HZ84, Pur88]. interaction [Hal87]. Interactions [Lan86]. interactivity [Pin80]. interdependence [Hai89]. interdisciplinarity [Ami85]. interface [Hal87]. International [BM83, IM89, McC80, SZB83, BD80, BS82, CCM88, EC87, Has81, MG86, ZSB82a, ZSB82b, de 80, Man80a]. Internationalization [Gua88]. internationally [SW83]. interpretation [Eto88, Vel87]. interrelation [PK87]. Intormation [LD87]. Introduction [HIM88, ZM80b]. invalidity [Bon80]. invention [BJ81, Yan82]. inventors [Lev86]. investigation [Mor82]. investigations [BS82, Hai80]. Invisible [Mul85, Chu85]. Irvine [Kea91]. Issues [RBBH89].

J [Ano87a, Fra87]. J. [Kea91]. Japan [Mor89b, Yam89]. Japanese [EM83b, Hai82a]. John [Ano85a, Ano85b, Ano87a, Ano87b, Ano88b, Ano89a, Ano89b, Gar85]. Journal [Bon85, RBBH89, AH85, Bal85b, BGS89d, Kre83b, Kre83a, Luu89, MT89, Sni84]. Journal-to-journal [RBBH89]. journals [AM88, BR85, Bal85c, Bal86, BN82, BGS89b, CN80, Dor85, Dor87, FP81, KAA^{+}88, Mid83, OV89a, PP85, SGB89, Sha87, Sla83, SE82, Tod83, Tod84, Yam89, Yan81, ZSB82a, ZSB82b, vRH87]. judgement [Koe82, MBFV85b]. judgments [PCJ88, Shr85].

Kingdom [CIM86, HIM88]. know [Pav88]. knowledge [Sen89c]. Kunz [Hai88].

laboratory [Coh80, Coh81, GR88, Qur84]. laboratory/group [Qur84]. Language [BJ83, BJ84, Swa88, SK86, Swa85, Yam89]. last [Tro87]. late [JNR86]. Latin [KPV86]. laude [NV87b, NV89]. law [Egg87, GS85, Han89, Bon80, CL87, Egg87, Eto84, Gup87, Gup89, KLC84, Kre83a, Kyv89, Sen89b, Nar89]. leaders [NG89]. learning [Kyv89]. Lectin [KS89b]. legitimate [CG80]. less [PABK88]. lesser [Fra80]. Lessons [Gie81, Mor89b]. Letter [KM88, Lin87a]. leximappe [TV89]. lib [Hje82]. library [Sit87]. Life [Hai84, HG86, Inh84, Mor84, Nal84, BGS87b, BGS88b, Dia84, Dia87, Ned85, SR89]. life-cycle [Dia84, Dia87]. Life-span [HG86]. Linear [Han89, Sen89b]. lines [KPV86]. literature [Ave85, BJ83, BJ84, BS82, BZ85, Chr85, DE89, Gup89, Hal89, Hur81, Hur84, KS89b, Mot89, Nor87, PL87, RMM87, SFL89, SK86, Sen85, Sen89c, Ste85, Tod85, VR88]. Little [CM85, Gar85, Lin88, PK87]. local [Gro89]. look [SG84]. Lotka [Egg87, Gup87, Gup89, Kre83a, Kun88b, Kyv89].

M [Kor89a]. M. [Hai88]. macroeconomics [McC83]. macrolevel [BS82]. main [Mot89]. Mainstream [DE89, AM88, Dav83, Oro81]. major [SGB89]. makers [DHQ89]. making [Por88]. man [Per86]. management [DS89, Fra84b]. managerial [NG89]. managers [Dan87]. manpower [Kuh80, Raj83, Col80]. manufacturing [DL85]. manuscript [Lin88]. Mapping [Kry89, Sma88, TV89, LBCW88, Ley87b, SSG85]. maps [KL85, RC84]. marginality [Shr85]. Martin [Kea91]. Mass [NG86]. mathematical [KB81, Sla83]. mathematics [BN82, BGS87a, BGS88a, PP85]. matrices [Nom82a, Sla83, TDV87, de 81a, de 81b]. matter [Tod83]. MBD [Van89b]. meaning [Vel86]. measurable [Lin89]. measure [ABT88, Dor85, Dor87, JSCL82, Kre83b, Lin89]. Measurement [Okr87, Ano84d, Ley89b, Moe89, Nas88, Sim84, vRH87]. measurements [Kid88]. measures [Bin81, LM82b, NR88, Pin80, PCJ88, SB81, SMKG86].

Measuring [Fra80, AM89, Lin89, Pok80]. mechanics [Hur81]. medical [BR85, Poi84, SZB85, Sha87]. Meeting [BM83]. Meetings [SZB83]. memorial [Ano85a]. memoriam [Ano84c]. mentor [LM85]. Merging [Rou89]. method [Bon85, NR88, Nom82a, Raj89, Sha89]. methodical [Lan86]. Methodological [Gor84, FHNR81, Per83a]. methodology [Mor85b]. Methods [DD82, vRH87, Ley87b, Ley89a, Nar89, Vin86]. Mexican [CdA89]. Mexico [JNR86, JHBH88]. mice [Lin87b]. Michael [Ano87a, Fra87]. microbiology [Sen89d]. microlevel [BS82]. microscope [VJL89]. mid [SZB85]. mid-size [SZB85]. Migration [Har86, vHvVLD83, Col88]. Milestone [CC87]. millstone [CC87]. mining [Hal87]. mobility [VD87a, VD87b]. model [Coz89, DL85, Dia84, Dia87, EM83b, HMM88, Hub81, KB81, Kot87, Ono88, SG84, Sen86b, Sim86b, SBN84, TA86, Vin87]. modelling [Kor89b]. models [Cou89]. Modern [BM86]. molecular [MT89]. money [Dia85]. Monte [Sim86a]. month [SKWK80]. month-by-month [SKWK80]. Moravcsik [Ano87a, Fra87]. much [KPV86]. Mullins [SH83]. multicomponent [LD86, LD87]. multidimensional [Gor84, Hai84, Inh84, Mor84, Nal84, TV89]. Multidimensionality [Fra84a]. multidisciplinary [BGS89b]. Multiple [Sim86a, Ste82, Dia85, Gor80, Hef81, Lin82, LM82b, LM82a, MvR86, POV86, SB86b]. multiple-authored [Dia85]. Multiples [Sim86b]. Multisectional [NG86]. Multivariate [SMKG86]. must [Mor87b].

Nalimov [Ano88b, Ano89b, Gra89a]. Narin [Ano89a, Van89a]. nation [BdSP81]. National [BGS89b, BN82, HS89, KM80, Ley89b, MVVL85, PP85, ST86, CF82]. nations [Ned88]. natural [Hai89, Mcg88]. nature [Hai83, KAA^{+}88, BGS89b]. negotiation [JKW87]. NERC [Mcg88]. Netherlands [Le 80, vHvVLD83]. nets [TA86]. network [Cou89, Kor89b]. networks [Dor85, Dor87, Mar81a, SBN84]. neuroscience [Sen89c]. newest [BGS88c, BGS88b, BGS88a, BGS89c]. News [Ano88a, Bal83, FB80a, FB80b, FB80c, Far80a, Far80b, Far81a, Far81b, Far81c, Far81d, Far81e, Far82a, Far82b, Far82c, Far82d, Far82e, Far83a, Far83b, Far83c, Far83d, Far83e, Far84a, Far84b, Far84c, Far84d, Far84e, Far85a, Far85b, Far85c, Far85d, Far86a, Far86b, Far86c, Far86d, Far86e, Far86f, Far87a, Far87b, Far87c, Far87d, Far87e, Far87f, Far88a, Far88b, Far88c, Far88d, Far88e, Far89b, Far89c, Far89a, Far89d, Far89e, Ste81]. niche [Pur89]. Nigeria [Gup87, Gup89]. Nigerian [Ada81, Ehi88]. nomination [GG88]. Non [PABK88, DE89, Hai83, Hai86a, Hai86b, Yab85]. non-additivity [Hai86a, Hai86b]. Non-citation [PABK88]. non-Gaussian [Hai83, Yab85]. Nordic [Luu89]. normal [She80]. note [ER88, Has81, Mal81, Sen89b, SB86b]. Notes [ZM80a]. nuclear [Hal87, Tro87]. Numbers [LR84, Ley89b]. objectives [Ano84d]. obliteration [Oro87]. Observations [MvR86, Sni84]. Obsolescence [Mar85]. ocean [CIM86]. oeuvre [MvR86]. old [Ada81, Her88]. One [BGS87c, BGS87b, BGS87a, Lib85]. ones [Per83a]. Online [Per86, Woo88]. Operationalizing [Van89b]. optimal [Dia87]. optimization [Gra89b]. optimum [Han89]. ordinary [NV87b]. ores [Hal87]. Organic [Gua86]. Organization [Lyo80, Poi84, Nas88]. organizational [Gro89]. organizations [Alo85, KM80]. origins [LR84]. Ortega [Mea87, Oro87, CC87, Law87, MM87, NV87a, Oro85, Sni86, Sni87]. other [BN82, vHvVLD83]. Output [Bin81, Van81, BM82, BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89c, CGH^{+}88, Gro89, Ned85, Okr87, POV86, SZB85, SB86a, SMKG86, SW83, SBN84, VJL89].

Paper [Bin81, Kun88b]. Papers [Ste82, Ave85, Bal85b, Gor80, Lan85, Per83a, RRJ83, SR89, Swa85]. paradigm [Gre84, Koc83, Vel87]. parameters [BdSP81, Sen86a]. Part [Hai86a, Hai86b, VD87a, VD87b]. particle [SKWK80]. past [IM85, MBFV85b]. Patent [Pav85, BJ81, BD87, Kun87, Mor89a]. patenting [SW83]. pattern [OV89a]. Patterns [Alo85, Kid89, RMM87, Ave85, BJ83, BJ84, BGS89b, Gup87, Har86, Hop84, JNR86, MC85, MT89, Ned88, SR89, ZSB82a, ZSB82b]. Peer [NV87b, KCB82, MBFV85b, SZTB84, SDH82]. [Gup89]. 20 [Egg86]. group [Qur84]. scientific [POV86]. Yugoslavia [PABK88]. Perceived [Ehi88]. perceptions [SK89]. performance [Col88, CIM86, IM85, IM89, Koe82, Ley89b, Luu89, MBFV85b, Moe89, MVVL85, NG89, NV87b, NV89, SB81, SGB87, SK89]. performances [Kre87]. period [Gup87]. periodic [KL85, She80]. Periodical [Ste85, BJ83, BJ84]. periodicals [Sen86a, Sen89d]. peripheral [DE89, Vel86]. peripheries [CA89]. periphery [AM88, AM89, Gua86, PS84]. perish [Luu89]. Personality [RMP83]. personnel [Mal81]. perspective [BD80, Nor87, de 80]. Ph.D.s [Har86]. pharmaceutical [Gua88]. phased [Nad81]. phenomenological [Sen84]. phenomenon [MvR86]. phosphides [CM81]. physical [Bow80, She80]. physicist [LvE81]. physicists [vHvVLD83]. Physics [PS84, Ste82, BN82, BGS87a, BGS88a, GR88, MGFA87, Mid83, Nad81, NV87b, PP85, SKWK80, Tod83, Tod85, Tro87]. Place [Sit87, Sen84]. planetary [Hal81]. Planning [NG86, NK88, PNN87]. plant [Nor87]. play [Lyo84]. plots [BS88]. pluralism [Gor84]. Poisson [Sim86b]. Policy [LBCW88, Ano82b, Ave80, Bro80, CIM86, HIM88, Mcg88, Mor80, Oro87, PT88, Por88, SMKG86, de 81a]. Polish [Bia87]. Politics [Cre88, Hol84]. polymer [LD86, LD87]. popularization [Kid88, Kid89]. population [Ric84]. position [KM80]. Possibilities [Pav85]. Possible [OV89b, Mor87a]. post [Chu85]. post-1972 [Chu85]. potential [Bia87, JHBH88, ST86]. Power [Nar89]. practices [PP85]. precision [Lin88]. preclinical [SR89]. precocity [HG86]. preface [Ano85c]. preferences [Lan86]. Preliminary [Swa85, Ada81, AG85]. present [CM81]. present-day [CM81]. presentation [Mor85a]. Price [Ano85a, Ano85b, Ano87a, Ano87b, Ano88b, Ano89a, Ano89b, Fra87, GS85, Gra89a, Mor85a, Mor87c, Van88, Van89a, Ano84c, Coz85, Egg87, Gar85, GS85, Gri84, Mar85, Moe89]. primary [Hai89]. printed [VJL89]. problem [Hur81, Van89b, Zuc87]. Problems [Hai86a, Hai86b, FHNR81, Mot89, Pav85, Pok80]. process [Kre87, Lin88, Mor82, SDH82]. processes [Pur88]. producer [Kea91]. production [Dav83, Gor80, Pok80]. Productivity [Kyv89, Dia84, Dia87, GR88, Gup87, Hub81, KM80, Kre85, Lea85, LM82b, MG86, Okr87, OV89a, RM81, Sen86b, SHC89, VD87a, VD87b, Yab80, Yan82]. professional [KM80, NG89]. professions [She80]. professors [RMP83]. profile [BM83]. profiles [CGH^{+}88]. programme [GG88]. programs [Ami85, Ned88]. progress [PCJ88]. projects [Ano82b]. Prolific [Lev86]. properties [SE82]. proposed [Sen89b]. prospects [IM85]. protein [CIM86]. psychological [LvE81]. psychology [SE82]. public [LC80]. Publication [Coh80, Coh81, KM80, Man80b, Qur84, ST86, BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89b, BGS89c, BGS89d, Cre88, Gor80, Has81, Ned88, Rou87, SB86a, Vin88c, Yan82]. publications [DH89, Gil89, Ley89b, SG83, Vin86, Vin88c, Vla85, vRH87]. Publish [Luu89]. published [SDH82]. publishing [MC85, SB81]. Pulp [Bin81]. purposes [Vin88a].

Qualitative [Cou89, PL86, Ley89a, SE82]. Quality [Shr85, Law86, LR84, Lin89]. Quantitative [DH89, Fra84b, HIM88, PK87, Sch83a, Sch83b, Sch83c, SZTB84, Sch84a, Sch84b, Sch84c, Sch85a, Sch85b, Sch86, Sch87, Sch88a, Sch88b, Sch89a, Sch89b, Ami85, AM89, Cou89, Dav83, Hai86a, Hai86b, Mcg88, Por88, Raj83, She82, Vin87]. quantum [Hur81]. Quasi [TDV87, Vin87]. Quasi-correspondence [TDV87]. quasi-quantitative [Vin87]. queries [ZM80a]. questions [Mor87b]. quick [NR88].

R&D [DS89, Eto84, Fre80, GG88, MVVL85, NG89, Pok80, Ros80, SHC89]. radiation [De 87]. radio [Gie81]. random [NR88]. rank [BS82, Hai83, Hub81, Mok89]. rank-frequency [Hub81]. ranking [Bon85, FHNR81]. rapid [SKWK80]. Rate [Hal89, BJ81, Coh80, Coh81, Qur84]. rates [Flo88, SB86b]. rather [Lin89]. ratio [Flo88]. Rationales [KLC84]. ray [Gie81]. re [Sni86]. re-examination [Sni86]. real [GG88]. real-time [GG88]. reality [Cre88]. reassessment [Gor80, Kea91]. recently [Sen89b]. recipient [Fra87, Gra89a, Mor87c, Van88, Van89a]. recipients [Sni84]. recognized [KCB82]. Recombinant [SG80]. records [Cre88]. Referativnyj [Ste85]. Referee [Yam89]. reference [Bal85c, SBN84]. references [Ley86, Mar81b, Sha89, Sma87, Sza85]. reflected [Ste85]. reflection [Kre83a]. Reflections [ER88]. regularities [Yab80]. reinterpretation [Mor87a]. related [Per83b]. relation [SMKG86]. relational [BGS89c, SB86a]. relations [Gor80, Ley89a]. relationship [Ave85, POV86]. relationships [VD87a, VD87b]. Relative [SB86a, BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89c, Vin86, Vin88c]. reliability [Plo89, RBBH89, SG83]. report [BS82]. Representation [Kre83b]. Representing [Tod89]. reprint [Swa88]. Reputation [Cre88]. request [Swa88]. rerank [Sen86a, Sen89d]. Research [Bia87, CF82, Col88, Mcg88, RM81, ABT88, Ami85, Ano82b, Bal85a, BCC89, Bin81, Bow80, CGH^{+}88, Coh80, Coh81, Cre88, CdA89, CIM86, Dan87, Dia84, Dia87, DT85, FN87, Gra89b, Gua86, Gua88, Gup87, Hal87, Hef81, Her88, JSCL82, JKW87, JNR86, Kin88, Koe82, Kor89b, Kre85, Kre87, Lan86, LBCW88, Law86, Lea85, LD86, Luu89, MC85, MvR86, Moe89, Mor89a, Mor87a, Mor89b, NK88, Ned85, Okr87, Per86, Pin80, Poi84, Por88, PC85, Pou89, PNN87, Rie80, RL85, RMP83, SR89, SB81, SZB85, SGB87, Sen86a, Sen86b, SMKG86, SK89, SG86, TK81, VJL89, Van81, Vel87]. researchers [RMM87]. resources [CF82]. restraint [Hol84]. resulting [EM83b]. results [Bon85, DD82, Mor87a, Vel87]. revenue [RM81]. review [BCC89, Lin88, NV87b, SZTB84, SDH82]. Reviews [KS89a, Bro88, BN84, Bru81, EM83a, GMS84, Gri81, HFG82, Inh81, IS82, Jac87, KDYR87, LM88, McA81, ML80, MM81, MPMG82, Mor83, MM85, Mor85c, Mor88, NM86, RFH^{+}83, RMD86, SM83, SMR84, SZTB84, YT89]. revised [Dor87]. revisited [CM85, IM89]. revolution [Gre83, Gre84]. revolutionary [She80]. rewarding [Lea85]. rhetoric [Coz89]. rhetoric-first [Coz89]. Rich [BZ85]. Rising [Eto88]. rock [Sni84]. role [KM80, NG89, Por88, Shr85, VR88]. roles [Per83b]. roll [Lin88]. root [GS85]. Roy [CC82]. Rule [Gup89, Egg86]. run [BD87].

S [JHBH88]. scales [SGB87]. scaling [Nom82b, TV89]. scatter [PL87, SK86, BS88]. scattering [Sen89a]. scholarly [Ove88, RM81]. scholarship [Woo88]. school [Mal89]. Science [Ave80, BD80, Bro80, DNW80, Hai89, KPV86, Mor80, Ada81, Ano82b, Ano84d, AG85, AM89, BM86, BGS89a, BGS89d, Chu85, CM85, CIM86, Dav83, DH89, Ehi88, EC87, FP81, FKM81, Gar85, Gie81, Gri84, Hal81, HIM88, Hol84, IM85, KAA^{+}88, Kea91, Kid88, Kid89, KL89, LC80, LS87, Ley87b, Ley89a, Ley89b, Lib85, Lin89, Man80a, Mar81a, McC80, Mcg88, MG89, Mor86a, Mul85, NN85, Nor87, PT88, Pur88, PK87, Raj83, Raj89, Sch83a, Sch83b, Sch83c, Sch84a, Sch84b, Sch84c, Sch85a, Sch85b, Sch86, Sch87, Sch88a, Sch88b, Sch89a, Sch89b, SGB89, She80, Sit87, SSG85, Sma88, SW83, Ste85, SBN84, TA86, ZSB82b, ZM80a, de 80, SS85, BGS89b]. science-policy [HIM88]. Sciences [Roc83, Bow80, BGS87b, BGS88b, Chr85, Coz85, CdA89, Hai89, Hal81, Lan85, Moe89, Nas88, NZDD89, Per83b, Rie80, SR89, TK81, Woo88]. Scientific [BM82, JNR86, NG86, SZB83, Sim84, AM88, BZ85, CN80, Col80, DE89, Fra80, FN87, FD86, GR88, Gor80, Gre83, Gre84, Hai83, Hai86a, Hai86b, Han89, HMM88, HZ84, HG85, Hub81, Hur84, IM89, Kea91, KCB82, Kot87, Kre87, Kuh80, LBCW88, Law86, Ley89b, Lyo80, Mal89, Mal81, Man80b, MG86, MGFA87, Mor82, NV87b, Okr87, Plo89, PCJ88, PABK88, Rou87, SG83, ST86, Sha81, SMKG86, Sim80, Swa88, Tod84, Tod89, VD87a, VD87b, VP89, Vin86, Vin88b, Vla85, Yab80, Yam89, Yan81, Yut86, vRH87]. scientifically [Vel86]. scientist [Bon80, HG85, Lea85, PCJ88]. Scientists [Nal87, Alo85, CCM88, Dia84, Dia87, FHNR81, HG86, LPP^{+}86, Lyo84, Ric84]. Scientometric [BS88, CM81, Gup89, Hai80, Mal89, SZB83, SZB85, SGB89, SE82, Tro87, Hai82b, Hai82c, Hai82d, Ley89a, Mok89, Nom82a, Pou89, SB81, SG84, Sha89, Ste87, TDV87, Vin88a, Vin88b, de 81a, de 81b]. Scientometrics [Gra89b, Sit87, Coh84, Mor85b, RC84, Sen84, Sza85, Yab85]. score [Vin88c]. scores [SGB87, SE82]. SDI [Bon85]. search [PABK88, Sni87]. second [SDH82]. section [Ste85]. sectors [Dav83]. selection [DL85]. selective [Bon85]. SEPSU [Col88]. serials [Hai82a]. Series [NG86, BJ81, Lyo80, Sim80]. sessions [Lyo80]. set [SGB89]. sets [Rou89]. seventies [JNR86]. Shifting [HG85]. Short [BD87]. Short-run [BD87]. shoulders [Lin87b]. sicence [Gar85]. significance [AL89, Sma87, Yan81]. significant [Her88]. signs [Rie80]. simulation [Raj89]. simulations [Sim86a]. simultaneous [Nom82b]. since [Kea91]. Singapore [AG85]. single [ABT88, Bon80, Dia85, SB86b]. single-authored [Dia85]. site [DL85]. Situation [vHvVLD83]. Six [MGFA87, CCM88, NK88, Ned88, PNN87]. size [Ada81, Coh80, Coh81, Kre85, Pav88, Qur84, SZB85]. skew [EM83b, SG84]. small [AG85, Ano88b, Van88]. social [Chu85, Gri84, HG85, Kre83b, LvE81, Mor82, NZDD89, Per83b, Pok80, Rie80, TK81]. social-psychological [LvE81]. socio [Ano84d, BS88, Pur89]. socio-cultural [Pur89]. socio-economic [Ano84d, BS88]. sociology [Oro81, Oro85, Per83a, Sni84]. Solla [Ano84c, Ano85a, Ano85b, Ano87a, Ano87b, Ano88b, Ano89a, Ano89b, Fra87, Gar85, Gra89a, Mor85a, Mor87c, Van88, Van89a]. Some [BGS89d, FHNR81, Kid88, Law86, LvE81, Pok80, Por88, SB81, Sim86a, SDH82, Sni84, BN82, Vin86, Vin88b, Yan82, BGS89a]. source [Sen89a]. sources [Hai82a, LPP^{+}86, Vla85]. South [Pou89]. Soviet [DH89, FP81, Yan82]. space [Nom82b]. span [HG86]. Spanish [MG86, MGFA87]. speakers [Lyo80]. special [SBN84]. specialization [Ric84]. specialties [Har86, Hur84, KB81]. spectra [Kun87]. square [GS85, Nom82a, de 81b]. Stable [Yab85]. standing [Dor85, Dor87]. state [Ehi88]. States [BJ81, BD87, DL85, LR84, RM81]. Stationary [Hai82b, Hai82c, Hai82d]. Statistical [Coh84, Nar89, Plo89, SG83, SMKG86]. statistics [Fra84a, Mok89, Pav85]. status [NG89]. stimulation [Ned88]. Stochastic [EM83b, Kot87, Ono88]. strategy [Gil89]. stratified [Dor85, Dor87]. strength [ST86]. strengths [MVVL85]. Structural [SBN84]. Structure [Ric84, Kre83b, Kre83a, Kre85, Kre87, LS87, McC83, Pav88, Sit87]. structures [TV89, VP89]. student [FKM81]. students [RMM87]. studies [Ano82b, Chu85, Col88, DH89, Gri84, Hai89, HIM88, Ley89a, Mcg88, Sch83a, Sch83b, Sch83c, Sch84a, Sch84b, Sch84c, Sch85a, Sch85b, Sch86, Sch87, Sch88a, Sch88b, Sch89a, Sch89b, Sen84, Ste87]. study [AG85, Ave85, CCM88, CM81, FKM81, GR88, Gre83, Gre84, Gup89, HS89, Hai88, Hal87, Kin88, Kun88a, LPP^{+}86, LKO80, Lib85, Lyo80, MBFV85b, NK88, NZDD89, NV89, POV86, PABK88, PNN87, RL85, She82]. Studying [KL89, DD82, Mot89]. subauthorship [Hef81]. subfield [VP89]. subfields [HS89, LD87, Mid83, SGB89, Vin88b]. Subject [SGB87, CN80, LKO80]. subjects [Bal85b]. Subsaharan [Dav83]. success [FKM81]. Sudden [Gre83, Gre84]. superconductivity [DD82, Nad81]. superseded [Gre84]. Supplement [Hje82]. supporting [Mar85]. survey [PS84]. surveying [Vin88a]. Switching [Le 80]. syndrome [Gre83, Gre84, SFL89]. synergetic [Pur88]. system [BS82, JSCL82, JKW87, VR88]. systems [LS87, LD86, LD87, Yam89].

T [JHBH88]. table [KL85]. tail [Eto88]. tails [Kun88b]. teaching [RMP83]. team [SHC89]. technical [Shr85]. technically [LR84]. techniques [AM89, Kin88]. Techno [Sim80]. Techno-scientific [Sim80]. technological [EM83b, Kea91, Mor86b, VJL89]. Technology [Mor86b, Ano82b, Ano84d, DL85, EC87, FP81, Fra84b, LC80, Ley89a, McC80, NN85, Pav88, Por88, Raj83, Raj89, Sma88, SW83, Man80a]. Teletraffic [KL89]. terms [Ley89b]. test [Hur81]. Testing [CC87, MM87, Sni87]. tests [Nar89]. Thailand [Yut86]. their [Ave85, Gil89, Gor80, Her88, LD87, Sen86a]. Theorem [Rou87]. theoretical [Per83a, Sen89b, SKWK80]. theory [Coh84, Coz85, Flo88, Gra89b, HMM88, Hop84, Ley87a, Ley89a, Moe89, Nad81, NV87a, Sha81, She80, SH83]. therefrom [Vin88b]. thin [DD82]. thin-film [DD82]. third [Fra87, Mor87c, BM82]. Three [Sen86a, Coh81, Hur84, MG89]. Tibor [Ano87b, Mor87c]. Time [Kun87, BJ81, GG88, Her88, Mar81b, MBFV85a, Sim80]. time-dependent [MBFV85a]. time-series [Sim80]. tip [SZTB84]. tool [CG80, Per86]. tools [Cou89, DHQ89]. topical [Kid89]. total [RRJ83]. Trade [Por88, VR88]. trained [LR84]. transaction [Nom82a, TDV87]. transactions [de 81b]. Transfer [NG86, Alo85]. Transition [BS82]. treatment [OV89b]. trends [BD87, LD87, Sza85]. tribute [Gar85]. Trita [Hje82]. Trita-lib-2013 [Hje82]. twentieth [Hal81, Her86, Her88]. Two [Col88, Bal85c, Vin88c]. type [Sen89b]. types [BGS89d].

U.S. [BD80, FP81, Har86, McC80, de 80]. UK [Cre88, PT88]. undergraduate [RMM87]. underlying [Mor81]. Understanding [SKWK80]. United [BJ81, BD87, CIM86, DL85, HIM88, LR84, RM81]. units [DT85, NK88, Okr87, SK89]. Universities [Ada81, Ehi88, Le 80, RM81]. university [FKM81, RM81, RMP83, Sit87]. upon [Pok80, Sla83]. USA [Mal81]. usage [HMM88]. Use [Ste87, BJ83, BJ84, BCC89, GG88, Kin88, RL85, SW83]. usefulness [NZDD89]. Using [Coz85, Lin89, Rou87, Mor89a, SS85, SSG85, SBN84]. USSR [Hai80]. utility [Has81]. utilization [Flo88].

V [Gra89a]. V. [Gra89a]. valid [Lin89]. validation [NV89]. validity [CC82, RBBH89, SH83]. value [Dia85, LR84]. values [BM86, Mok89]. variables [Hai89]. Various [Ley87b]. Vasilevich [Ano88b, Ano89b]. Vasilij [Ano88b, Ano89b]. Velho [Mor87a]. verification [EM83b]. version [BGS87c, BGS87b, BGS87a, BGS88c, BGS88b, BGS88a, BGS89c, Nar89]. versus [BS88, Nor87]. veterinary [RMM87]. virology [Kin88]. visible [Luu89, SZTB84]. Vital [Rie80].

Wanner [Sim86b]. war [Sim80]. warfare [BJ81]. Warsaw [Sit87]. wastage [Flo88]. water [HS89]. weaknesses [MVVL85]. weightage [Sen89d]. Weighted [Vin88c]. weights [Tod84]. well [Fra84a]. were [Her88]. where [BR85]. Which [Hai89]. whins [Ano89b]. Who [BR85]. win [Ano88b]. Winning [Lyo84]. wins [Ano85b, Ano87a, Ano87b, Ano89a]. Women [CCM88]. word [LBCW88, RC84, TV89]. work [CCM88, HZ84, Ned85, SK89]. working [Ave85]. World [Ano82b, Mor86b, AG85, CN80, Gor84, Hai84, HG85, Inh84, Mor80, Mor84, Nal84, PS84, BM82].

X [Gie81]. X-ray [Gie81].

Yablonsky [BIM^{+}87]. yearly [Sim80]. years [MGFA87, Tro87].

Zealand [Bal85a]. Zhurnal [Ste85]. Zipf [Kun88b].

References

Ajiferuke:1988:CCS

[ABT88] Isola Ajiferuke, Q. Burell, and Jean Tague. Collaborative coefficient: A single measure of the degree of collaboration in research. Scientometrics, 14(5–6):421–433, November 1988. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

           http://link.springer.com/article/10.1007/BF02017100.

Adamson:1981:SSO

[Ada81] I. Adamson. The size of science in the old Nigerian universities — a preliminary analysis. Scientometrics, 3(4):317–324, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

           2861 (electronic). URL http://link.springer.com/article/
           10.1007/BF02021124.

Arunachalam:1985:SCW

[AG85] S. Arunachalam and K. C. Garg. A small country in a world of big science a preliminary bibliometric study of science in Singapore. Scientometrics, 8(5–6):301–313, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02018055.

Arunachalam:1985:JAA

[AH85] S. Arunachalam and S. Hiranaiah. Has journal of astrophysics and astronomy a future? Scientometrics, 8(1–2):3–11, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02025218.

Amsterdamska:1989:CIS

[AL89] Olga Amsterdamska and L. Leydesdorff. Citations: Indicators of significance? Scientometrics, 15(5–6):449–471, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017065.

Aloni:1985:PIT

[Alo85] Michaela Aloni. Patterns of information transfer among engineers and applied scientists in complex organizations. Scientometrics, 8(5–6):279–300, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018054.

Arunachalam:1988:HDJ

[AM88] S. Arunachalam and K. Manorama. How do journals on the periphery compare with mainstream scientific journals? Scientometrics, 14(1–2):83–95, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020244.

Arunachalam:1989:CBQ

[AM89] S. Arunachalam and K. Manorama. Are citation-based quantitative techniques adequate for measuring science on the periphery? Scientometrics, 15(5–6):393–408, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017061.

Amir:1985:DIR

[Ami85] S. Amir. On the degree of interdisciplinarity of research programs: A quantitative assessment. Scientometrics, 8(1–2):117– 136, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025225.

Anonymous:1981:CBa

[Ano81a] Anonymous. Call for bibliographies. Scientometrics, 3(1): 3–4, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02021859.

Anonymous:1981:CBb

[Ano81b] Anonymous. Call for bibliographies. Scientometrics, 3(3):157–158, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02101662.

Anonymous:1981:CBc

[Ano81c] Anonymous. Call for bibliographies. Scientometrics, 3(4):263–264, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02021120.

Anonymous:1981:CBd

[Ano81d] Anonymous. Call for bibliographies. Scientometrics, 3(5):347– 348, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02017572.

Anonymous:1981:CBe

[Ano81e] Anonymous. Call for bibliographies. Scientometrics, 3(6):413– 414, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02017434.

Anonymous:1982:EAB

[Ano82a] Anonymous. Editorial and advisory board. Scientometrics, 4(1):3, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02098002.

Anonymous:1982:WDR

[Ano82b] Anonymous. World directory of research projects, studies and courses in science and technology policy. Scientometrics, 4(6): 457–462, November 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021141.

Anonymous:1983:E

[Ano83]  Anonymous.  Editorial.  Scientometrics, 5(2):83, March

1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02072854.

Anonymous:1984:A

[Ano84a] Anonymous. Announcement. Scientometrics, 6(2):147, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02021290.

Anonymous:1984:C

[Ano84b] Anonymous. Correction. Scientometrics, 6(5):367, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/pdf/
10.1007/BF02020133.pdf.

Anonymous:1984:MDS

[Ano84c] Anonymous. In memoriam: Derek de Solla Price. Scientometrics, 6(1):3, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02020107.

Anonymous:1984:IMI

[Ano84d] Anonymous. Indicators of measurement of impact of science and technology on socio-economic development objectives. Scientometrics, 6(6):449–463, November 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02025832.

Anonymous:1985:DJS

[Ano85a] Anonymous. Derek John de Solla Price memorial award. Scientometrics, 7(3–6):139, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02017139.

Anonymous:1985:EGW

[Ano85b] Anonymous. Eugene Garfield wins the 1984 Derek John de Solla Price Award. Scientometrics, 7(3–6):141, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017140.

Anonymous:1985:IP

[Ano85c] Anonymous. Instead of a preface. Scientometrics, 7(3– 6):137, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02017138.

Anonymous:1987:MJM

[Ano87a] Anonymous. Michael J. Moravcsik wins the 1985 Derek John de Solla Price Award. Scientometrics, 11(3–4):123, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016585. See comments [Fra87].

Anonymous:1987:TBW

[Ano87b] Anonymous. Tibor Braun wins the 1986 Derek John de Solla Price Award. Scientometrics, 11(5–6):261, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/BF02279347.

See comments [Mor87c].

Anonymous:1988:N

[Ano88a] Anonymous. News. Scientometrics, 13(1–2):71–80, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02095765.

Anonymous:1988:VVN

[Ano88b] Anonymous. Vasilij Vasilevich Nalimov and Henry Small win the 1987 Derek John de Solla Price Award. Scientometrics, 14(5– 6):359, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02017095. See comment

[Van88].

Anonymous:1989:FNW

[Ano89a] Anonymous. Francis Narin wins the 1988 Derek John de Solla Price Award. Scientometrics, 17(1–2):3, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/BF02017716.

Anonymous:1989:VVN

[Ano89b] Anonymous. Vasilij Vasilevich Nalimov whins the 1987 Derek John de Solla Price Award. Scientometrics, 15(1–2):5, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02021793. See comment [Gra89a].

Averch:1980:SIP

[Ave80] H. Averch. Science indicators and policy analysis. Scientometrics, 2(5–6):339–345, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02095076.

Aversa:1985:CPH

[Ave85] Elizabeth Smith Aversa. Citation patterns of highly cited papers and their relationship to literature aging: A study of the working literature. Scientometrics, 7(3–6):383–389, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017156.

Balazs:1983:N

[Bal83] Judit Balázs. News. Scientometrics, 5(2):139–142, March 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02072860.

Balog:1985:ARN

[Bal85a] Catherine Balog. Agricultural research in New Zealand. Scientometrics, 8(1–2):81–89, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02025222.

Balog:1985:APD

[Bal85b] Catherine Balog. Authorship of papers dealing with different subjects in an agricultural journal. Scientometrics, 7(1–2):105–109, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020144.

Balog:1985:DRC

[Bal85c] Catherine Balog. The distribution of reference citations in two agricultural journals. Scientometrics, 7(1–2):101–104, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020143.

Balog:1986:IFG

[Bal86] Catherine Balog. Information flow to genetics journals. Scientometrics, 9(1–2):51–57, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016607.

Balaban:1987:AA

[Bal87] A. T. Balaban. Aurel Avramescu (1903–1985). Scientometrics, 11(1–2):3–5, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016623.

Bastide:1989:URA

[BCC89] F. Bastide, J. P. Courtial, and M. Callon. The use of review articles in the analysis of a research area. Scientometrics, 15 (5–6):535–562, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017070.

Ben-David:1980:USI

[BD80] J. Ben-David. “U.S. science in international perspective”. Scientometrics, 2(5–6):411–421, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095084. See com-

ments [de 80].

Brunk:1987:SRT

[BD87] G. G. Brunk and G. Demack. Short-run trends in United States patent activity. Scientometrics, 12(1–2):111–133, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016693.

Burke:1981:DCN

[BdSP81] Clare E. Burke and D. de S. Price. The distribution of citations from nation to nation on a field by field basis — a computer calculation of the parameters. Scientometrics, 3(5):363–377, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017574.

Braun:1987:OMVc

[BGS87a] T. Braun, W. Glänzel, and A. Schubert. One more version of the facts and figures on publication output and relative citation impact in physics and mathematics 1978–1980. Scientometrics, 12(1–2):3–16, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016686.

Braun:1987:OMVb

[BGS87b] T. Braun, W. Glänzel, and A. Schubert. One more version of the facts and figures on publication output and relative citation impact in the life sciences and chemistry 1978–1980. Scientometrics, 11(3–4):127–140, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016587.

Braun:1987:OMVa

[BGS87c] T. Braun, W. Glänzel, and A. Schubert. One more version of the facts and figures on publication output and relative citation impact of 107 countries 1978–1980. Scientometrics, 11(1–2):9–15, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016625.

Braun:1988:NVFc

[BGS88a] T. Braun, W. Glänzel, and A. Schubert. The newest version of the facts and figures on publication output and relative citation impact in physics, engineering and mathematics 1981–1985. Scientometrics, 14(5–6):365–382, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017097.

Braun:1988:NVFb

[BGS88b] T. Braun, W. Glänzel, and A. Schubert. The newest version of the facts and figures on publication output and relative citation impact in the life sciences and chemistry 1981–1985. Scientometrics, 14(1–2):3–15, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020240.

Braun:1988:NVFa

[BGS88c] T. Braun, W. Glänzel, and A. Schubert. The newest version of the facts and figures on publication output and relative citation impact of 100 countries 1981–1985. Scientometrics, 13(5–6):181– 188, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02019957.

Braun:1989:AAB

[BGS89a] T. Braun, W. Glänzel, and A. Schubert. Assessing assessments of British science. Some facts and figures to accept or decline. Scientometrics, 15(3–4):165–170, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017195.

Braun:1989:NPP

[BGS89b] T. Braun, W. Glänzel, and A. Schubert. National publication patterns and citation impact in the multidisciplinary journals Nature and Science. Scientometrics, 17(1–2):11–14, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017719.

Braun:1989:NVF

[BGS89c] T. Braun, W. Glänzel, and A. Schubert. The newest version of the facts and figures on publication output and relative citation impact. A collection of relational charts, 1981–1985. Scientometrics, 15(1–2):13–20, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021795.

Braun:1989:SDD

[BGS89d] T. Braun, W. Glänzel, and A. Schubert. Some data on the distribution of journal publication types in the science citation index database. Scientometrics, 15(5–6):325–330, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017057.

Bialon:1987:RDP

[Bia87] Lidia Bialoñ. Research and development potential of Polish industry. Scientometrics, 12(3–4):163–177, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016290.

Blauberg:1987:Y

[BIM^{+}87] Igor V. Blauberg, Andrey A. Ignatyev, Eduard M. Mirsky, Vadim N. Sadovsky, and Alexander P. Uzdemir. A. I. Yablonsky (1936–1986). Scientometrics, 11(1–2):7–8, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016624.

Bindon:1981:OMC

[Bin81] G. Bindon. Output measures of cooperative research: The case of the Pulp and Paper Research Institute of Canada. Scientometrics, 3(2):85–106, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02025632.

Brunk:1981:IWR

[BJ81] G. G. Brunk and G. J. Jason. The impact of warfare on the rate of invention: A time series analysis of United States patent activity. Scientometrics, 3(6):437–455, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017436.

Baldauf:1983:LUP

[BJ83] R. B. Baldauf and B. H. Jernudd, Jr. Language use patterns in the fisheries periodical literature. Scientometrics, 5(4):245–255, July 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02019740.

Baldauf:1984:LUP

[BJ84] R. B. Baldauf, Jr. and B. H. Jernudd. Language use patterns in the fisheries periodical literature. Scientometrics, 6(1):67–69, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/content/
pdf/10.1007/BF02020116.pdf.

Blickenstaff:1982:C

[Bli82] J. Blickenstaff. Correction. Scientometrics, 4(4):345, July 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). See [BM82].

Blickenstaff:1982:SOT

[BM82] J. Blickenstaff and M. J. Moravcsik. Scientific output in the Third World. Scientometrics, 4(2):135–169, March 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02018451. See correction [Bli82].

Blickenstaff:1983:PIM

[BM83] J. Blickenstaff and M. J. Moravcsik. The profile of an international meeting. Scientometrics, 5(3):143–154, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02095624.

Blanpied:1986:MSH

[BM86] W. A. Blanpied and M. J. Moravcsik. Modern science and human values. Scientometrics, 10(1–2):119–123, July 1986. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016864.

Braun:1982:CES

[BN82] T. Braun and J. I. Nagy. A comparative evaluation of some Hungarian and other national biology, chemistry, mathematics and physics journals. Scientometrics, 4(6):439–455, November 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021140.

Brown:1984:BR

[BN84] W. B. Brown and V. V. Nalimov. Book reviews. Scientometrics, 6(3):203–205, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016764.

Bonitz:1980:EIB

[Bon80] M. Bonitz. Evidence for the invalidity of the Bradford Law for the single scientist. Scientometrics, 2(3):203–214, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016698.

Bonitz:1985:JRS

[Bon85] M. Bonitz. Journal ranking by selective impact. New method based on SDI results and journal impact factors. Scientometrics, 7(3–6):471–485, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017162.

Bowers:1980:IBR

[Bow80] R. Bowers. Indicators of basic research in the physical sciences. Scientometrics, 2(5–6):429–433, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02095086.

Bakker:1985:EMJ

[BR85] P. Bakker and H. Rigter. Editors of medical journals: Who and from where. Scientometrics, 7(1–2):11–22, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020137.

Braun:1989:E

[Bra89] T. Braun. Editorial. Scientometrics, 17(3–4):193–194, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02026408.

Brooks:1980:SIS

[Bro80] H. Brooks. Science indicators and science policy. Scientometrics, 2(5–6):331–337, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02095075.

Broadus:1987:TDB

[Bro87] R. N. Broadus. Toward a definition of “bibliometrics”. Scientometrics, 12(5–6):373–379, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016680.

Bromley:1988:BR

[Bro88] D. A. Bromley. Book reviews. Scientometrics, 13(5–6):301–303, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02019966.

Bruer:1981:BR

[Bru81] John T. Bruer. Book reviews. Scientometrics, 3(4):335–337, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021126.

Bonitz:1982:TMM

[BS82] M. Bonitz and P. Schmidt. Transition from the macrolevel to the microlevel of information at rank distribution investigations of the report literature of an international information system. Scientometrics, 4(4):283–295, July 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02021644.

Braun:1988:SVS

[BS88] T. Braun and A. Schubert. Scientometric versus socio-economic indicators. Scatter plots for 51 countries. 1978–1980. Scientometrics, 13(1–2):3–9, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02095758.

Braun:1985:GSL

[BZ85] T. Braun and S. Zsindely. Growth of scientific literature and the barnaby Rich effect. Scientometrics, 7(3–6):529–530, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017165.

Chatelin:1989:BCP

[CA89] Y. Chatelin and R. Arvanitis. Between centers and peripheries. Scientometrics, 17(5–6):437–452, November 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017464.

Campbell:1982:VRA

[CC82] C. A. Campbell and M. B. M. Campbell. The validity of the Roy approximation for citation analysis. Scientometrics, 4(6): 411–416, November 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021137.

Cole:1987:TOH

[CC87] S. Cole and J. R. Cole. Testing the Ortega Hypothesis: Milestone or millstone? Scientometrics, 12(5–6):345–353, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016678.

Chakravarthy:1988:WSW

[CCM88] R. Chakravarthy, A. Chawla, and G. Mehta. Women scientists at work — an international comparative study of six countries. Scientometrics, 14(1–2):43–74, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02020242.

Cronin:1989:GDM

[CdA89] Blaise Cronin and Judith Licea de Arenas. The geographic distribution of Mexican health sciences research. Scientometrics, 17(1–2):39–48, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017721.

Chapman:1982:CRN

[CF82] I. D. Chapman and C. Farina. Concentration of resources: The National Research Council’s (Canada) grants in aid of research: 1964–1974. Scientometrics, 4(2):105–117, March 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02018449.

Chubin:1980:CAL

[CG80] D. Chubin and E. Garfield. Is citation analysis a legitimate evaluation tool? Scientometrics, 2(1):91–94, January 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016602.

Carpenter:1988:BPB

[CGH^{+}88] M. P. Carpenter, F. Gibb, M. Harris, J. Irvine, B. R. Martin, and F. Narin. Bibliometric profiles for British academic institutions: An experiment to develop research output indicators. Scientometrics, 14(3–4):213–233, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02020076.

Christovao:1985:ALB

[Chr85] H. T. Christovão. The aging of the literature of biomedical sciences in developed and developing countries. Scientometrics, 7 (3–6):411–430, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017158.

Chubin:1985:BIC

[Chu85] D. E. Chubin. Beyond invisible colleges: Inspirations and aspirations of post-1972 social studies of science. Scientometrics, 7 (3–6):221–254, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017148.

Crouch:1986:BAS

[CIM86] D. Crouch, J. Irvine, and B. R. Martin. Bibliometric analysis for science policy: An evaluation of the United Kingdom’s research performance in ocean currents and protein crystallography. Scientometrics, 9(5–6):239–267, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017247.

Chen:1987:BLI

[CL87] Ye-Sho Chen and F. F. Leimkuhler. Bradford’s Law: An index approach. Scientometrics, 11(3–4):183–198, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016591.

Chernogorenko:1981:SEP

[CM81] V. B. Chernogorenko and S. V. Muchnik. Scientometric estimation of present-day study on phosphides. Scientometrics, 3(2):127–134, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025635.

Cole:1985:LSB

[CM85] S. Cole and G. S. Meyer. Little science, big science revisited. Scientometrics, 7(3–6):443–458, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017160.

Carpenter:1980:SCW

[CN80] M. P. Carpenter and F. Narin. The subject composition of the world’s scientific journals. Scientometrics, 2(1):53–63, January 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016599.

Cohen:1980:PRF

[Coh80] J. E. Cohen. Publication rate as a function of laboratory size in a biomedical research institution. Scientometrics, 2(1):35–52, January 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016598.

Cohen:1981:PRF

[Coh81] J. E. Cohen. Publication rate as a function of laboratory size in three biomedical research institutions. Scientometrics, 3(6): 467–487, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017438.

Cohen:1984:STA

[Coh84] J. E. Cohen. Statistical theory aids inference in scientometrics. Scientometrics, 6(1):27–32, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02020111.

Cole:1980:CIS

[Col80] S. Cole. Comments on “Indicators of scientific manpower”. Scientometrics, 2(5–6):405–409, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095083. See [Kuh80].

Collins:1988:RPM

[Col88] P. M. D. Collins. Research performance and migration: Two SEPSU studies. Scientometrics, 14(3–4):201–211, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020075.

Courtial:1989:QMQ

[Cou89] J. P. Courtial. Qualitative models, quantitative tools and network analysis. Scientometrics, 15(5–6):527–534, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017069.

Cozzens:1985:UAD

[Coz85] Susan E. Cozzens. Using the archive: Derek Price’s theory of differences among the sciences. Scientometrics, 7(3–6):431–441, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017159.

Cozzens:1989:WDC

[Coz89] Susan E. Cozzens. What do citations count? the rhetoric-first model. Scientometrics, 15(5–6):437–447, May 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017064.

Crewe:1988:RRR

[Cre88] I. Crewe. Reputation, research and reality: The publication records of UK departments of politics, 1978–1984. Scientometrics, 14(3–4):235–250, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020077.

Daniels:1987:CII

[Dan87] W. D. Daniels. Choosing input indicators for research managers. Scientometrics, 11(1–2):17–25, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016626.

Davis:1983:ISM

[Dav83] C. H. Davis. Institutional sectors of ‘mainstream’ science production in Subsaharan Africa, 1970–1979: A quantitative analysis. Scientometrics, 5(3):163–175, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095626.

Dobrov:1982:MRS

[DD82] G. M. Dobrov and I. V. Dziekovskaya. Methods and results of studying the flow of information in the field of thin-film superconductivity. Scientometrics, 4(1):27–44, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02098004.

deSollaPrice:1980:CUS

[de 80] D. de Solla Price. Comments on “U.S. science in an international perspective”. Scientometrics, 2(5–6):423–428, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095085. See [BD80].

deSollaPrice:1981:ASMa

[de 81a] D. de Solla Price. The analysis of scientometric matrices for policy implications. Scientometrics, 3(1):47–53, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021863.

deSollaPrice:1981:ASMb

[de 81b] D. de Solla Price. The analysis of square matrices of scientometric transactions. Scientometrics, 3(1):55–63, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021864.

DeStefano:1987:CAA

[De 87] D. A. De Stefano. Citation analysis and adaptive radiation. Scientometrics, 11(1–2):43–51, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016628.

Davis:1989:MNM

[DE89] Ch. H. Davis and Th. O. Eisemon. Mainstream and non mainstream scientific literature in four peripheral Asian scientific communities. Scientometrics, 15(3–4):215–239, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017200.

Doroshenko:1989:QSS

[DH89] S. I. Doroshenko and S. D. Haitun. Quantitative studies of science: a bibliography of Soviet publications. Scientometrics, 15(1– 2):139–154, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021805.

Dou:1989:IAT

[DHQ89] H. Dou, P. Hassanaly, and L. Quoniam. Infographic analytical tools for decision makers. Scientometrics, 17(1–2):61–70, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017723.

Diamond:1984:EML

[Dia84] A. M. Diamond, Jr. An economic model of the life-cycle research productivity of scientists. Scientometrics, 6(3):189–196, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016762.

Diamond:1985:MVC

[Dia85] A. M. Diamond, Jr. The money value of citations to singleauthored and multiple-authored articles. Scientometrics, 8(5–6): 315–320, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018056.

Diamond:1987:OCM

[Dia87] A. M. Diamond, Jr. An optimal control model of the life-cycle research productivity of scientists. Scientometrics, 11(3–4):251–253, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016595.

Davies:1985:SSM

[DL85] Barbara R. Davies and J. Martin Lazniarz. A site selection model for high technology manufacturing firms in the United States. Scientometrics, 8(1–2):103–116, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02025224.

Dewitt:1980:SCI

[DNW80] T. W. Dewitt, R. S. Nicholson, and M. K. Wilson. Science citation index and chemistry. Scientometrics, 2(4):265–275, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016348.

Doreian:1985:MSJ

[Dor85] P. Doreian. A measure of standing of journals in stratified networks. Scientometrics, 8(5–6):341–363, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02018058.

Doreian:1987:RMS

[Dor87] P. Doreian. A revised measure of standing of journals in stratified networks. Scientometrics, 11(1–2):71–80, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016631.

Dobrov:1989:FEA

[DS89] G. Dobrov and A. Skofenko. Fuzzy expertise and its application to R&D management. Scientometrics, 15(1–2):21–31, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021796.

Dobrov:1985:CAE

[DT85] G. M. Dobrov and V. E. Tonkal. Comparative analysis and estimation of competence of research units. Scientometrics, 7(3– 6):309–325, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017151.

Eto:1987:ABD

[EC87] H. Eto and P. M. Candelaria. Applicability of the Bradford distribution to international science and technology indicators. Scientometrics, 11(1–2):27–42, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016627.

Egghe:1986:R

[Egg86] L. Egghe. On the 80/20 rule. Scientometrics, 10(1–2):55–68, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016860.

Egghe:1987:ECP

[Egg87] L. Egghe. An exact calculation of Price’s Law for the law of Lotka. Scientometrics, 11(1–2):81–97, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016632.

Ehikhamenor:1988:PSS

[Ehi88] F. A. Ehikhamenor. Perceived state of science in Nigerian universities. Scientometrics, 13(5–6):225–238, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02019960.

Eisemon:1983:BR

[EM83a] Thomas Eisemon and Michael J. Moravcsik. Book reviews. Scientometrics, 5(6):405–407, November 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02134327.

Eto:1983:SMI

[EM83b] H. Eto and K. Makino. Stochastic model for innovation and resulting skew distribution for technological concentration with verification in Japanese industry. Scientometrics, 5(4):219–243, July 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02019739.

Egghe:1988:RDN

[ER88] L. Egghe and R. Rousseau. Reflections on a deflection: A note on different causes of the Groos droop. Scientometrics, 14(5–6): 493–511, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017105.

Eto:1984:BLR

[Eto84] H. Eto. Bradford Law in R&D expending of firms and R&D concentration. Scientometrics, 6(3):183–188, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016761.

Eto:1988:RTB

[Eto88] H. Eto. Rising tail in Bradford distribution: Its interpretation and application. Scientometrics, 13(5–6):271–287, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02019963.

Farkas:1980:Nd

[Far80a] J. Farkas. News. Scientometrics, 2(4):321–323, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016352.

Farkas:1980:Ne

[Far80b] J. Farkas. News. Scientometrics, 2(5–6):451–461, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02095090.

Farkas:1981:Na

[Far81a] J. Farkas. News. Scientometrics, 3(1):67, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/BF02021866.

Farkas:1981:Nb

[Far81b] J. Farkas. News. Scientometrics, 3(2):149–154, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02025637.

Farkas:1981:Nc

[Far81c] J. Farkas. News. Scientometrics, 3(4):339–343, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021127.

Farkas:1981:Nd

[Far81d] J. Farkas. News. Scientometrics, 3(5):403–410, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017579.

Farkas:1981:Ne

[Far81e] J. Farkas. News. Scientometrics, 3(6):493–498, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017440.

Farkas:1982:Na

[Far82a] J. Farkas. News. Scientometrics, 4(1):83–86, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02098009.

Farkas:1982:Nb

[Far82b] J. Farkas. News. Scientometrics, 4(2):175–179, March 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02018453.

Farkas:1982:Nc

[Far82c] J. Farkas. News. Scientometrics, 4(3):235–239, May 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021063.

Farkas:1982:Nd

[Far82d] J. Farkas. News. Scientometrics, 4(5):405–410, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02135127.

Farkas:1982:Ne

[Far82e] J. Farkas. News. Scientometrics, 4(6):463–467, November 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021142.

Farkas:1983:Na

[Far83a] J. Farkas. News. Scientometrics, 5(1):75–82, January 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02097179.

Farkas:1983:Nb

[Far83b] J. Farkas. News. Scientometrics, 5(3):205–209, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02095630.

Farkas:1983:Nc

[Far83c] J. Farkas. News. Scientometrics, 5(4):273–276, July 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02019743.

Farkas:1983:Nd

[Far83d] J. Farkas. News. Scientometrics, 5(5):335–342, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02147230.

Farkas:1983:Ne

[Far83e] J. Farkas. News. Scientometrics, 5(6):409, November 1983. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/10.1007/
BF02134328.

Farkas:1984:Na

[Far84a] J. Farkas. News. Scientometrics, 6(1):63–66, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02020115.

Farkas:1984:Nb

[Far84b] J. Farkas. News. Scientometrics, 6(2):139–146, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021289.

Farkas:1984:Nc

[Far84c] J. Farkas. News. Scientometrics, 6(3):207–213, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016765.

Farkas:1984:Nd

[Far84d] J. Farkas. News. Scientometrics, 6(4):267–278, July 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02279361.

Farkas:1984:Ne

[Far84e] J. Farkas. News. Scientometrics, 6(5):371–379, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02020135.

Farkas:1985:Na

[Far85a] J. Farkas. News. Scientometrics, 7(1–2):123–131, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02020147.

Farkas:1985:Nb

[Far85b] J. Farkas. News. Scientometrics, 8(1–2):145–148, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02025228.

Farkas:1985:Nc

[Far85c] J. Farkas. News. Scientometrics, 8(3–4):263–275, September 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016940.

Farkas:1985:Nd

[Far85d] J. Farkas. News. Scientometrics, 8(5–6):377–382, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02018060.

Farkas:1986:Na

[Far86a] J. Farkas. News. Scientometrics, 9(1–2):99–101, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016611.

Farkas:1986:Nb

[Far86b] J. Farkas. News. Scientometrics, 9(3–4):193–196, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017242.

Farkas:1986:Nc

[Far86c] J. Farkas. News. Scientometrics, 9(5–6):307–313, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017252.

Farkas:1986:Nd

[Far86d] J. Farkas. News. Scientometrics, 10(1–2):125–132, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016865.

Farkas:1986:Ne

[Far86e] J. Farkas. News. Scientometrics, 10(3–4):227–234, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02026044.

Farkas:1986:Nf

[Far86f] J. Farkas. News. Scientometrics, 10(5–6):329–333, November 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016778.

Farkas:1987:Na

[Far87a] J. Farkas. News. Scientometrics, 11(1–2):115–122, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016635.

Farkas:1987:Nb

[Far87b] J. Farkas. News. Scientometrics, 11(3–4):255–259, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016596.

Farkas:1987:Nc

[Far87c] J. Farkas. News. Scientometrics, 11(5–6):367–369, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02279355.

Farkas:1987:Nd

[Far87d] J. Farkas. News. Scientometrics, 12(1–2):145–148, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016695.

Farkas:1987:Ne

[Far87e] J. Farkas. News. Scientometrics, 12(3–4):259–265, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016296.

Farkas:1987:Nf

[Far87f] J. Farkas. News. Scientometrics, 12(5–6):421–424, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016685.

Farkas:1988:Na

[Far88a] J. Farkas. News. Scientometrics, 13(3–4):177–179, March 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017183.

Farkas:1988:Nb

[Far88b] J. Farkas. News. Scientometrics, 13(5–6):305–310, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02019967.

Farkas:1988:Nc

[Far88c] J. Farkas. News. Scientometrics, 14(1–2):169–175, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02020251.

Farkas:1988:Nd

[Far88d] J. Farkas. News. Scientometrics, 14(3–4):351–357, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020085.

Farkas:1988:Ne

[Far88e] J. Farkas. News. Scientometrics, 14(5–6):521–527, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017107.

Farkas:1989:Nc

[Far89a] J. Farkas. News. Scientometrics, 15(1–2):159–163, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021807.

Farkas:1989:Na

[Far89b] J. Farkas. News. Scientometrics, 15(3–4):317–323, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017208.

Farkas:1989:Nb

[Far89c] J. Farkas. News. Scientometrics, 15(5–6):633–638, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017075.

Farkas:1989:Nd

[Far89d] J. Farkas. News. Scientometrics, 17(1–2):183–190, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017733.

Farkas:1989:Ne

[Far89e] J. Farkas. News. Scientometrics, 17(3–4):381–385, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02026420.

Farkas:1980:Na

[FB80a] J. Farkas and J. Balázs. News. Scientometrics, 2(1):85–90, January 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016601.

Farkas:1980:Nb

[FB80b] J. Farkas and Judith Balázs. News. Scientometrics, 2(2):169–176, March 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02279366.

Farkas:1980:Nc

[FB80c] J. Farkas and Judith Balázs. News. Scientometrics, 2(3):243–249, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016702.

Frijdal:1986:CAS

[FD86] A. Frijdal and J. P. De Greve. Communication activities in scientific disciplines in Belgium. Scientometrics, 9(1–2):37–49, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016606.

Folly:1981:SMP

[FHNR81] G. Folly, B. Hajtman, J. I. Nagy, and I. Ruff. Some methodological problems in ranking scientists by citation analysis. Scientometrics, 3(2):135–147, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02025636.

Frieze:1981:AUS

[FKM81] Irene Hanson Frieze, J. M. Knoble, and I. I. Mitroff. American university students’ beliefs about success in science: A case study. Scientometrics, 3(2):115–126, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02025634.

Flor:1988:IWR

[Flo88] A. G. Flor. The information wastage ratio: Towards a theory of rates of information generation and utilization. Scientometrics, 13 (1–2):63–69, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095764.

Frame:1987:GCS

[FN87] J. Davidson Frame and F. Narin. The growth of Chinese scientific research, 1973–84. Scientometrics, 12(1–2):135–144, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016694.

Frame:1981:CUS

[FP81] J. D. Frame and D. R. Prokrym. Counts of U.S. and Soviet science and technology journals. Scientometrics, 3(3):159–175, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02101663.

Frey:1988:ADA

[FP88] B. S. Frey and W. W. Pommerehne. The American domination among eminent economists. Scientometrics, 14(1–2):97–110, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020245.

Frame:1980:MSA

[Fra80] J. D. Frame. Measuring scientific activity in lesser developed countries. Scientometrics, 2(2):133–145, March 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02279364.

Frame:1984:MAW

[Fra84a] J. D. Frame. Multidimensionality is alive and well in applied statistics. Scientometrics, 6(2):97–101, March 1984. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021283.

Frame:1984:QMT

[Fra84b] J. D. Frame. Quantitative management of technology. Scientometrics, 6(4):223–232, July 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02279357.

Frame:1987:CMJ

[Fra87] J. Davidson Frame. Comments on Michael J. Moravcsik, recipient of the third Derek de Solla Price Award. Scientometrics, 11(3– 4):125–126, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02016586. See [Ano87a].

Freeman:1980:IIR

[Fre80] R. B. Freeman. Indicators of the impact of R&D on the economy. Scientometrics, 2(5–6):375–385, October 1980. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02095080. See comments [Ros80].

Garfield:1985:TDJ

[Gar85] E. Garfield. In tribute to Derek John de Solla Price: a citation analysis of little science, big sicence. Scientometrics, 7(3– 6):487–503, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017163.

Giusti:1988:UCN

[GG88] W. L. Giusti and L. Georghiou. The use of co-nomination analysis in real-time evaluation of an R&D programme. Scientometrics, 14(3–4):265–281, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020079.

Gieryn:1981:ASE

[Gie81] T. F. Gieryn. The aging of a science and its exploitation of innovation: Lessons from X-ray and radio astronomy. Scientometrics, 3(4):325–334, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021125.

Gillett:1989:DBD

[Gil89] R. Gillett. Determining the best departments by their best publications: A strategy best avoided. Scientometrics, 17(1–2):121– 125, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017728.

Gieryn:1984:BR

[GMS84] Thomas F. Gieryn, Michael J. Moravscik, and Frederick Suppe. Book reviews. Scientometrics, 6(4):261–266, July 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02279360.

Gordon:1980:CRI

[Gor80] M. D. Gordon. A critical reassessment of inferred relations between multiple authorship, scientific collaboration, the production of papers and their acceptance for publication. Scientometrics, 2(3):193–201, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016697.

Gordon:1984:MPM

[Gor84] M. D. Gordon. Methodological pluralism in a multidimensional world. Scientometrics, 6(2):87–92, March 1984. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021281.

Garg:1988:BAS

[GR88] K. C. Garg and M. K. D. Rao. Bibliometric analysis of scientific productivity: A case study of an Indian physics laboratory. Scientometrics, 13(5–6):261–269, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02019962.

Granovsky:1989:CVV

[Gra89a] Yu V. Granovsky. Comments on V. V. Nalimov recipient of the 1987 Derek de Solla Price Award. Scientometrics, 15(1–2):7–12, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021794. See [Ano89b].

Granovsky:1989:STE

[Gra89b] Yu V. Granovsky. Scientometrics, theory of experiment and optimization of research. Scientometrics, 15(1–2):33–43, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021797.

Gregory:1983:CSS

[Gre83] J. G. Gregory. Citation study of a scientific revolution: Sudden infant death syndrome. Scientometrics, 5(5):313–327, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02147227.

Gregory:1984:CSS

[Gre84] J. G. Gregory. Citation study of a scientific revolution: Sudden infant death syndrome 2. The superseded paradigm. Scientometrics, 6(5):307–326, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020130.

Griffith:1981:BR

[Gri81] Belver C. Griffith. Book reviews. Scientometrics, 3(3):251–252, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02101669.

Griffith:1984:DPS

[Gri84] B. C. Griffith. Derek Price (1922–1983) and the social studies of science. Scientometrics, 6(1):5–7, January 1984. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020108.

Groenewegen:1989:ILO

[Gro89] P. Groenewegen. Influences of local and organizational factors on output indicators. Scientometrics, 15(5–6):409–422, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017062.

Glanzel:1985:PDE

[GS85] W. Glänzel and A. Schubert. Price distribution. An exact formulation of Price’s “square root law”. Scientometrics, 7(3–6):211–219, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017147.

Guay:1986:EBR

[Gua86] Y. Guay. Emergence of basic research on the periphery: Organic chemistry in India, 1907–1926. Scientometrics, 10(1–2):77–94, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016862.

Guay:1988:IIR

[Gua88] Y. Guay. Internationalization of industrial research: The pharmaceutical industry; 1965–1979. Scientometrics, 13(5–6):189–213, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02019958.

Gupta:1987:LLP

[Gup87] D. K. Gupta. Lotka’s Law and productivity patterns of entomological research in Nigeria for the period, 1900–1973. Scientometrics, 12(1–2):33–46, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016688.

Gupta:1989:SSB

[Gup89] D. K. Gupta. Scientometric study of biochemical literature of Nigeria, 1970–1984: Application of Lotka’s Law and the 80 /20rule. Scientometrics, 15(3–4):171–179, March 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017196.

Haitun:1980:SIU

[Hai80] S. D. Haitun. Scientometric investigations in the USSR. Scientometrics, 2(1):65–84, January 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016600.

Haigh:1982:CAF

[Hai82a] M. J. Haigh. Citation analysis of foreign sources in Japanese geographical serials. Scientometrics, 4(3):195–203, May 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021060.

Haitun:1982:SSDa

[Hai82b] S. D. Haitun. Stationary scientometric distributions. Scientometrics, 4(1):5–25, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02098003.

Haitun:1982:SSDb

[Hai82c] S. D. Haitun. Stationary scientometric distributions. Scientometrics, 4(2):89–104, March 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018448.

Haitun:1982:SSDc

[Hai82d] S. D. Haitun. Stationary scientometric distributions. Scientometrics, 4(3):181–194, May 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021059.

Haitun:1983:RDE

[Hai83] S. D. Haitun. The “rank distortion” effect and non-Gaussian nature of scientific activities. Scientometrics, 5(6):375–395, November 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02134325.

Haitun:1984:LMW

[Hai84] S. D. Haitun. Life in a multidimensional world. Scientometrics, 6(2):93–96, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021282.

Haitun:1986:PQAa

[Hai86a] S. D. Haitun. Problems of quantitative analysis of scientific activities: The non-additivity of data. Part I. Scientometrics, 10 (1–2):3–16, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016856.

Haitun:1986:PQAb

[Hai86b] S. D. Haitun. Problems of quantitative analysis of scientific activities: The non-additivity of data. Part II. Corollaries. Scientometrics, 10(3–4):133–155, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02026038.

Haitun:1988:MKA

[Hai88] S. D. Haitun. On M. Kunz’s article “A case study against Haitun’s conjectures”. Scientometrics, 13(1–2):35–44, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095761.

Haitun:1989:SSN

[Hai89] S. D. Haitun. Science studies and natural sciences: Which is primary, distribution or interdependence between variables? Scientometrics, 15(1–2):45–58, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02021798.

Hall:1981:EPS

[Hal81] D. H. Hall. The earth and planetary sciences in science during the twentieth century. Scientometrics, 3(5):349–362, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017573.

Hall:1987:IBG

[Hal87] D. H. Hall. The interface between geoscience and industry: A case study of the interaction between research and the discovery and mining of ores for nuclear fuels. Scientometrics, 11(3–4):199–216, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016592.

Hall:1989:RGL

[Hal89] D. H. Hall. Rate of growth of literature in geoscience from computerized databases. Scientometrics, 17(1–2):15–38, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017720.

Han:1989:LIL

[Han89] Haichao Han. Linear increase law of optimum age of scientific creativity. Scientometrics, 15(3–4):309–312, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017206.

Hargens:1986:MPU

[Har86] L. L. Hargens. Migration patterns of U.S. Ph.D.s among disciplines and specialties. Scientometrics, 9(3–4):145–164, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017238.

Hastings:1981:NUI

[Has81] T. Hastings. A note on the utility of international publication data. Scientometrics, 3(5):389–396, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017576.

Heffner:1981:FRM

[Hef81] A. G. Heffner. Funded research, multiple authorship, and subauthorship collaboration in four disciplines. Scientometrics, 3 (1):5–12, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021860.

Herrmann:1986:ATC

[Her86] D. B. Herrmann. Astronomy in the twentieth century. Scientometrics, 9(3–4):187–191, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017241.

Herrmann:1988:HOW

[Her88] D. B. Herrmann. How old were the authors of significant research in twentieth century astronomy at the time of their greatest achievements? Scientometrics, 13(3–4):135–137, March 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017180.

Haitun:1982:BR

[HFG82] S. D. Haitun, J. Davidson Frame, and Belver C. Griffith. Book reviews. Scientometrics, 4(1):77–81, January 1982. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02098008.

Hongzhou:1985:SWS

[HG85] Zhao Hongzhou and Jiang Guohua. Shifting of world’s scientific center and scientists’ social ages. Scientometrics, 8(1–2):59–80, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025221.

Hongzhou:1986:LSP

[HG86] Zhao Hongzhou and Jiang Gouhua. Life-span and precocity of scientists. Scientometrics, 9(1–2):27–36, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016605.

Healey:1988:IQS

[HIM88] P. Healey, J. Irvine, and B. R. Martin. Introduction: Quantitative science-policy studies in the United Kingdom. Scientometrics, 14(3–4):177–183, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02020073.

Hjerppe:1982:SBB

[Hje82] R. Hjerppe. Supplement to a “bibliography of bibliometrics and citation indexing & analysis” (trita-lib-2013). Scientometrics, 4 (3):241–273, May 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021064.

Hohenester:1988:UTM

[HMM88] A. Hohenester, L. Mathelitsch, and M. J. Moravcsik. The usage of ‘theory’ and ‘model’ in scientific conceptualization. Scientometrics, 14(5–6):411–420, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017099.

Holmfeld:1984:PRS

[Hol84] J. D. Holmfeld. Politics and the restraint of science. Scientometrics, 6(5):369–370, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02020134.

Hopkins:1984:NCT

[Hop84] Frances L. Hopkins. New causal theory and ethnomethodology: Cocitation patterns across a decade. Scientometrics, 6(1):33–53, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020112.

Hagendijk:1989:ANS

[HS89] R. P. Hagendijk and J. W. Smeenk. The analysis of national subfields: a case study of Dutch fresh-water ecology. Scientometrics, 15(5–6):485–508, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017067.

Hubert:1981:RFM

[Hub81] J. J. Hubert. A rank-frequency model for scientific productivity. Scientometrics, 3(3):191–202, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02101665.

Hurt:1981:TDL

[Hur81] C. D. Hurt. A test of differences in the literature history of four historical accounts of the quantum mechanics problem. Scientometrics, 3(6):457–466, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017437.

Hurt:1984:ELD

[Hur84] C. D. Hurt. An examination of the literature distributions of three scientific specialties. Scientometrics, 6(2):115–126, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021287.

Hong-Zhou:1984:ICS

[HZ84] Zhao Hong-Zhou. An intelligence constant of scientific work. Scientometrics, 6(1):9–17, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020109.

Irvine:1985:EBS

[IM85] J. Irvine and B. R. Martin. Evaluating big science: CERN’s past performance and future prospects. Scientometrics, 7(3– 6):281–308, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017150.

Irvine:1989:ICS

[IM89] J. Irvine and B. R. Martin. International comparisons of scientific performance revisited. Scientometrics, 15(5–6):369–392, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017060. See reply [Kea91].

Inhaber:1981:BR

[Inh81] H. Inhaber. Book reviews. Scientometrics, 3(5):401–402, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017578.

Inhaber:1984:LMW

[Inh84] H. Inhaber. Life in a multidimensional world. Scientometrics, 6 (2):105–107, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021285.

Inhaber:1982:BR

[IS82] H. Inhaber and A. Schubert. Book reviews. Scientometrics, 4 (2):171–173, March 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018452.

Jaccerino:1987:BR

[Jac87] V. Jaccerino. Book reviews. Scientometrics, 12(5–6):413– 417, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016683.

Jimenez:1988:PMH

[JHBH88] J. Jiménez, P. Hunya, M. Bayona, and A. Halász. The s & t potential of Mexico and Hungary. Scientometrics, 14(1–2):17–41, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020241.

Jakubowski:1987:ARF

[JKW87] A. Jakubowski, R. Kulikowski, and D. Wagner. Allocation of research funds in competitive environment. A computerized negotiation system. Scientometrics, 12(3–4):179–196, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016291.

Jimenez:1986:SRA

[JNR86] J. Jiménez, M. A. Navarro, and M. W. Rees. Scientific research areas in Mexico: Growth patterns in the late seventies. Scientometrics, 9(5–6):209–221, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017244.

Jagodzinski-Sigogneau:1982:HMD

[JSCL82] M. Jagodzinski-Sigogneau, J.-P. Courtial, and B. Latour. How to measure the degree of independence of a research system? Scientometrics, 4(2):119–133, March 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018450.

Kaneiwa:1988:CBJ

[KAA^{+}88] K. Kaneiwa, J. Adachi, M. Aoki, T. Masuda, A. Midorikawa, A. Tanimura, and S. Yamazaki. A comparison between the journals nature and science. Scientometrics, 13(3–4):125–133, March 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017179.

Kochen:1981:MGM

[KB81] M. Kochen and A. Blaivas. A model for the growth of mathematical specialties. Scientometrics, 3(4):265–273, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021121.

Kochen:1982:DSE

[KCB82] M. Kochen, R. Crickman, and A. Blaivas. Distribution of scientific experts as recognized by peer consensus. Scientometrics, 4(1):45–56, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02098005.

Kranzberg:1987:BR

[KDYR87] Melvin Kranzberg, David Eli Drew, Yongyuth Yuthavong, and Marcel Roche. Book reviews. Scientometrics, 11(1–2):105–113, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016634.

Kealey:1991:GFA

[Kea91] T. Kealey. Government-funded academic science is a consumer good, not a producer good: A comparative reassessment of Britain’s scientific and technological achievements since 1794 and a comment on the bibliometry of B. Martin and J. Irvine. Scientometrics, 20(2):369–394, February 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017527. See [IM89].

Kidd:1988:PSS

[Kid88] J. S. Kidd. The popularization of science: Some basic measurements. Scientometrics, 14(1–2):127–142, July 1988. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020247.

Kidd:1989:PSI

[Kid89] J. S. Kidd. The popularization of science II: Patterns of topical coverage. Scientometrics, 15(3–4):241–255, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017201.

King:1988:UBT

[Kin88] J. King. The use of bibliometric techniques for institutional research evaluation: A study of avian virology research. Scientometrics, 14(3–4):295–313, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02020081.

Kochen:1985:MDD

[KL85] M. Kochen and J. Lansing. On maps for discovery: Did the periodic table guide elemental discovery? Scientometrics, 7(3– 6):327–339, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017152.

Kranakis:1989:TCS

[KL89] E. Kranakis and L. Leydesdorff. Teletraffic conferences: Studying a field of engineering science. Scientometrics, 15(5–6):563–591, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017071.

Karmeshu:1984:RBL

[KLC84] Karmeshu, N. C. Lind, and V. Cano. Rationales for Bradford’s Law. Scientometrics, 6(4):233–241, July 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02279358.

Knorr:1980:PPP

[KM80] K. D. Knorr and R. Mittermeir. Publication productivity and professional position: Cross-national evidence on the role of organizations. Scientometrics, 2(2):95–120, March 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02279362.

Kruytbosch:1988:LE

[KM88] C. Kruytbosch and M. J. Moravcsik. Letter to editor. Scientometrics, 13(5–6):299–300, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02019965.

Kochen:1983:CP

[Koc83] Manfred Kochen. The cognitive paradigm. Scientometrics, 5(5): 331–333, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02147229.

Koenig:1982:DEJ

[Koe82] M. E. D. Koenig. Determinants of expert judgement of research performance. Scientometrics, 4(5):361–378, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02135122.

Korennoi:1989:GMD

[Kor89a] A. Korennoi. Gennady M. Dobrov 1929–1989. Scientometrics, 17(1–2):9–10, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02017718.

Korennoi:1989:ICM

[Kor89b] A. A. Korennoi. Information co-modelling of a network of research institutions. Scientometrics, 15(1–2):59–71, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021799.

Kot:1987:SME

[Kot87] S. M. Kot. The stochastic model of evolution of scientific disciplines. Scientometrics, 12(3–4):197–205, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016292.

Krauskopf:1986:SLA

[KPV86] M. Krauskopf, R. Pessot, and R. Vicuña. Science in Latin America how much and along what lines? Scientometrics, 10(3–4): 199–206, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026041.

Kretschmer:1983:RLL

[Kre83a] Hildrun Kretschmer. The reflection of Lotka’s Law in the structure of citations of a journal. Scientometrics, 5(2):85–92, March 1983.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02072855.

Kretschmer:1983:RCS

[Kre83b] Hildrun Kretschmer. Representation of a complex structure measure for social groups and its application to the structure of citations in a journal. Scientometrics, 5(1):5–30, January 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02097175.

Kretschmer:1985:CSG

[Kre85] Hildrun Kretschmer. Cooperation structure, group size and productivity in research groups. Scientometrics, 7(1–2):39–53, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020140.

Kretschmer:1987:ACS

[Kre87] Hildrun Kretschmer. The adaptation of the cooperation structure to the research process and scientific performances in research groups. Scientometrics, 12(5–6):355–372, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016679.

Kryzhanovsky:1988:ABH

[Kry88] L. N. Kryzhanovsky. An application of bibliometrics to the history of electricity. Scientometrics, 14(5–6):487–492, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017104.

Kryzhanovsky:1989:MHE

[Kry89] L. N. Kryzhanovsky. Mapping the history of electricity. Scientometrics, 17(1–2):165–170, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017730.

Kostyuk:1989:BR

[KS89a] V. Kostyuk and J. Schreider. Book reviews. Scientometrics, 15 (1–2):155–157, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021806.

Kumari:1989:GLL

[KS89b] L. Kumari and I. N. Sengupta. Growth of lectin literature: 1954– 1982. Scientometrics, 17(3–4):353–361, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02026417.

Kuh:1980:ISM

[Kuh80] Charlotte V. Kuh. Indicators of scientific manpower. Scientometrics, 2(5–6):395–403, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095082. See com-

ments [Col80].

Kunz:1987:TSP

[Kun87] M. Kunz. Time spectra of patent information. Scientometrics, 11 (3–4):163–173, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016589.

Kunz:1988:CSA

[Kun88a] M. Kunz. A case study against Haitun’s conjectures. Scientometrics, 13(1–2):25–33, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02095760.

Kunz:1988:LZP

[Kun88b] M. Kunz. Lotka and Zipf: Paper dragons with fuzzy tails. Scientometrics, 13(5–6):289–297, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019964.

Kyvik:1989:PDF

[Kyv89] S. Kyvik. Productivity differences fields of learning, and Lotka’s Law. Scientometrics, 15(3–4):205–214, March 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017199.

Lange:1985:EDC

[Lan85] Lydia Lange. Effects of disciplines and countries on citation habits. An analysis of empirical papers in behavioural sciences.

Scientometrics, 8(3–4):205–215, September 1985. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016936.

Lange:1986:IBD

[Lan86] Lydia Lange. Interactions between disciplines and countries in methodical preferences for empirical research. Scientometrics, 10(5– 6):281–295, November 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016775.

Lawani:1986:SBC

[Law86] S. M. Lawani. Some bibliometric correlates of quality in scientific research. Scientometrics, 9(1–2):13–25, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016604.

Lawani:1987:OHI

[Law87] S. M. Lawani. The Ortega Hypothesis, individual differences, and cumulative advantage. Scientometrics, 12(5–6):321–323, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016673.

Law:1988:PMS

[LBCW88] J. Law, S. Bauin, J.-P. Courtial, and J. Whittaker. Policy and the mapping of scientific change: A co-word analysis of research into environmental acidification. Scientometrics, 14(3–4): 251–264, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020078.

LaPorte:1980:IPA

[LC80] T. R. La Porte and D. Chisholm. Indicators of public attitudes toward science and technology. Scientometrics, 2(5–6):439–448, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095088.

Lipatov:1986:BIF

[LD86] Yu. S. Lipatov and L. V. Denisenko. On the behaviour of information flows in multicomponent polymer systems research. Scientometrics, 9(5–6):197–207, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017243.

Lipatov:1987:IFS

[LD87] Yu. S. Lipatov and L. V. Denisenko. Intormation flows in the subfields of multicomponent polymer systems and trends of their development. Scientometrics, 11(3–4):175–182, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016590.

LePair:1980:SBA

[Le 80] C. Le Pair. Switching between academic disciplines in universities in the Netherlands. Scientometrics, 2(3):177–191, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016696.

Leary:1985:FAR

[Lea85] R. A. Leary. A framework for assessing and rewarding a scientist’s research productivity. Scientometrics, 7(1–2):29–38, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020139.

Levine:1986:PIB

[Lev86] L. O. Levine. Prolific inventors — a bibliometric analysis. Scientometrics, 10(1–2):35–42, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016858.

Leydesdorff:1986:DFR

[Ley86] L. Leydesdorff. The development of frames of references. Scientometrics, 9(3–4):103–125, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017235.

Leydesdorff:1987:TTC

[Ley87a] L. Leydesdorff. Towards a theory of citation? Scientometrics, 12(5–6):305–309, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016669.

Leydesdorff:1987:VMM

[Ley87b] L. Leydesdorff. Various methods for the mapping of science. Scientometrics, 11(5–6):295–324, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02279351.

Leydesdorff:1989:RBQ

[Ley89a] L. Leydesdorff. The relations between qualitative theory and scientometric methods in science and technology studies. Scientometrics, 15(5–6):333–347, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017058.

Leydesdorff:1989:SCI

[Ley89b] L. Leydesdorff. The science citation index and the measurement of national performance in terms of numbers of scientific publications. Scientometrics, 17(1–2):111–120, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017727.

Libkind:1985:OAS

[Lib85] A. N. Libkind. One approach to study communication in science. Scientometrics, 8(3–4):217–231, September 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016937.

Lindsey:1982:FEA

[Lin82] D. Lindsey. Further evidence for adjusting for multiple authorship. Scientometrics, 4(5):389–395, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02135124.

Line:1987:LE

[Lin87a] M. B. Line. Letter to editor. Scientometrics, 12(5–6):419, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016684.

Line:1987:SGB

[Lin87b] M. B. Line. The shoulders of giants, or the backs of mice? Scientometrics, 12(5–6):297–298, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF02016666.

Lindsey:1988:APM

[Lin88] D. Lindsey. Assessing precision in the manuscript review process: A little better than a dice roll. Scientometrics, 14(1–2):75–82, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020243.

Lindsey:1989:UCC

[Lin89] D. Lindsey. Using citation counts as a measure of quality in science measuring what’s measurable rather than what’s valid. Scientometrics, 15(3–4):189–203, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017198.

Lawson:1980:BSN

[LKO80] Julia Lawson, Barbara Kostrewski, and C. Oppenheim. A bibliometric study on a new subject field: Energy analysis. Scientometrics, 2(3):227–237, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016700.

Long:1982:FEA

[LM82a] J. S. Long and R. Mcginnis. Further evidence for adjusting for multiple authorship. Scientometrics, 4(5):397–398, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02135125.

Long:1982:APM

[LM82b] J. S. Long and R. McGinnis. On adjusting productivity measures for multiple authorship. Scientometrics, 4(5):379–387, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02135123.

Long:1985:EMA

[LM85] J. S. Long and R. McGinnis. The effects of the mentor on the academic career. Scientometrics, 7(3–6):255–280, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017149.

Lyon:1988:BR

[LM88] W. S. Lyon and R. Martin. Book reviews. Scientometrics, 13(3– 4):173–175, March 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017182.

Lancaster:1986:FIS

[LPP^{+}86] F. W. Lancaster, Maria A. Porta, Kathryn Plagenz, Krys Szymborski, and Marjorie Krebs. Factors influencing sources cited by scientists: A case study for Cuba. Scientometrics, 10(5–6): 243–257, November 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016773.

Lerner:1984:NOE

[LR84] J. Lerner and R. Roy. Numbers, origins, economic value and quality of technically trained immigrants into the United States. Scientometrics, 6(4):243–259, July 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02279359.

Lewicka-Strzalecka:1987:DSS

[LS87] Anna Lewicka-Strzalecka. Dynamics and structure of systems science. Scientometrics, 12(3–4):207–219, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016293.

Luukkonen:1989:PVJ

[Luu89] Terttu Luukkonen. Publish in a visible journal of perish? Assessing citation performance of Nordic cancer research. Scientometrics, 15(5–6):349–367, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017059.

Leydesdorff:1981:SSP

[LvE81] L. Leydesdorff and H. van Erkelens. Some social-psychological aspects of becoming a physicist. Scientometrics, 3(1):27–45, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021862.

Lyon:1980:OAS

[Lyo80] W. S. Lyon. Organization, attendance, speakers, and sessions: A study of four scientific conference series. Scientometrics, 2(3):215– 226, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016699.

Lyon:1984:WGS

[Lyo84] W. S. Lyon. Winning the games scientists play. Scientometrics, 6(1):61–62, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02020114.

Malecki:1981:NGC

[Mal81] E. J. Malecki. A note on the geographical concentration of scientific personnel in the USA. Scientometrics, 3(2):107–114, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025633.

Malciene:1989:SAS

[Mal89] L. Malciené. Scientometric analysis of a scientific school. Scientometrics, 15(1–2):73–85, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02021800.

Mansfield:1980:CII

[Man80a] E. Mansfield. Comments on “International indicators of science and technology”. Scientometrics, 2(5–6):369–373, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF02095079. See [McC80].

Manten:1980:PSI

[Man80b] A. A. Manten. Publication of scientific information is not identical with communication. Scientometrics, 2(4):303–308, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016350.

Marshakova:1981:CNI

[Mar81a] Irina V. Marshakova. Citation networks in information science. Scientometrics, 3(1):13–25, January 1981. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02021861.

Marton:1981:CTD

[Mar81b] J. Marton. Changes in the time distribution of biochemical article references from 1962 to 1977. Scientometrics, 3(5):397– 400, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017577.

Marton:1985:OIE

[Mar85] J. Marton. Obsolescence or immediacy? Evidence supporting Price’s hypothesis. Scientometrics, 7(3–6):145–153, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017142.

Moed:1985:ABI

[MBFV85a] H. F. Moed, W. J. M. Burger, J. G. Frankfort, and A. F. J. Van Raan. The application of bibliometric indicators: Important fieldand time-dependent factors to be considered. Scientometrics, 8(3– 4):177–203, September 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016935.

Moed:1985:CSB

[MBFV85b] H. F. Moed, W. J. M. Burger, J. G. Frankfort, and A. F. J. Van Raan. A comparative study of bibliometric past performance analysis and peer judgement. Scientometrics, 8(3–4):149–159, September 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016933.

McAllister:1985:EAB

[MC85] P. R. McAllister and T. Condon. Econometric analysis of biomedical research publishing patterns. Scientometrics, 7(1–2):55–75, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020141.

McAllister:1981:BR

[McA81] P. R. McAllister. Book reviews. Scientometrics, 3(1):65–66, January 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02021865.

McCulloch:1980:IIS

[McC80] Rachel McCulloch. International indicators of science and technology: How does the U.S. compare? Scientometrics, 2(5–6): 355–367, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095078. See comments [Man80a].

McCain:1983:ACS

[McC83] Katherine W. McCain. The author cocitation structure of macroeconomics. Scientometrics, 5(5):277–289, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02147224.

Mcginnety:1988:NER

[Mcg88] J. A. Mcginnety. The Natural Environment Research Council (NERC): Recent experiences with quantitative science policy studies. Scientometrics, 14(3–4):283–293, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020080.

Meadows:1987:OH

[Mea87] A. J. Meadows. Ortega hypothesis. Scientometrics, 12(5–6): 315–316, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02016671.

Mendez:1986:SSP

[MG86] A. Méndez and I. Gómez. The Spanish scientific productivity through eight international databases. Scientometrics, 10(3–4): 207–219, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026042.

Mendez:1989:CCC

[MG89] Aida Méndez and Isabel Gómez. A comparison of citation classics in three fields of science. Scientometrics, 15(5–6):621–631, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017074.

Mendez:1987:SYS

[MGFA87] A. Méndez, I. Gómez, M. T. Fernández, and G. López Aguado. Six years of Spanish scientific activity in physics and engineering through INSPEC and COMPENDEX. Scientometrics, 12 (1–2):81–100, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016691.

Midorikawa:1983:CAP

[Mid83] N. Midorikawa. Citation analysis of physics journals: Comparison of subfields of physics. Scientometrics, 5(6):361–374, November 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02134324.

Moravcsik:1980:BR

[ML80] M. J. Moravcsik and W. S. Lyon. Book reviews. Scientometrics, 2(3):239–242, May 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016701.

Moravcsik:1981:BR

[MM81] Edith A. Moravcsik and Joshua Menkes. Book reviews. Scientometrics, 3(6):489–492, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017439.

Moravcsik:1985:BRa

[MM85] M. J. Moravcsik and Jack Medows. Book reviews. Scientometrics, 7(1–2):119–121, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020146.

Macroberts:1987:TOH

[MM87] M. H. Macroberts and Barbara R. Macroberts. Testing the Ortega Hypothesis: Facts and artifacts. Scientometrics, 12(5–6): 293–295, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016665.

Moed:1989:BMR

[Moe89] H. F. Moed. Bibliometric measurement of research performance and Price’s theory of differences among the sciences. Scientometrics, 15(5–6):473–483, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017066.

Mokhov:1989:ASE

[Mok89] O. I. Mokhov. About statistics of the extreme values and the rank form of scientometric distributions. Scientometrics, 15(1– 2):87–96, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021801.

Moravcsik:1980:SSP

[Mor80] M. J. Moravcsik. Science and science policy in the Arab world. Scientometrics, 2(5–6):449–450, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF02095089.

Morse:1981:UCB

[Mor81] P. M. Morse. The underlying characteristics of the Bradford distribution. Scientometrics, 3(6):415–436, November 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017435.

Moravcsik:1982:SPS

[Mor82] Michael J. Moravcsik. The social process of scientific investigation. Scientometrics, 4(3):233–234, May 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF02021062.

Moravcsik:1983:BR

[Mor83] Michael J. Moravcsik. Book reviews. Scientometrics, 5(4):269– 272, July 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02019742.

Moravcsik:1984:LMW

[Mor84] M. J. Moravcsik. Life in a multidimensional world. Scientometrics, 6(2):75–85, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021280.

Moravcsik:1985:APF

[Mor85a] M. J. Moravcsik. Address at the presentation of the first Derek de Solla Price Award to Eugene Garfield on December 20, 1984. Scientometrics, 7(3–6):143–144, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017141.

Moravcsik:1985:ASA

[Mor85b] M. J. Moravcsik. Applied scientometrics: An assessment methodology for developing countries. Scientometrics, 7(3–6):165–176, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017144.

Moravcsik:1985:BRb

[Mor85c] M. J. Moravcsik. Book reviews. Scientometrics, 8(1–2):141–143, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025227.

Moravcsik:1986:CSS

[Mor86a] M. J. Moravcsik. The classification of science and the science of classification. Scientometrics, 10(3–4):179–197, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02026040.

Morgan:1986:TFD

[Mor86b] Robert P. Morgan. Technology, finance and development: An analysis of the world bank as a technological institution. Scientometrics, 9(5–6):305–306, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02017251.

Moravcsik:1987:BEP

[Mor87a] M. J. Moravcsik. In the Beholder’s eye: A possible reinterpretation of Velho’s results on Brazilian agricultural research. Scientometrics, 11(1–2):53–57, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016629.

Moravcsik:1987:WMA

[Mor87b] M. J. Moravcsik. We must ask questions before giving answers. Scientometrics, 12(5–6):299–301, November 1987. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016667.

Moravscik:1987:CTB

[Mor87c] M. J. Moravscik. Comments on Tibor Braun, recipient of the third Derek de Solla Price Award. Scientometrics, 11(5– 6):263–264, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02279348. See [Ano87b].

Moravcsik:1988:BR

[Mor88] Michael J. Moravcsik. Book reviews. Scientometrics, 14(3–4): 347–350, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020084.

Moral:1989:EDA

[Mor89a] L. P. Moral. Elements for a diagnosis of applied research and development in Cuba using patent information: 1968–1983. Scientometrics, 17(1–2):83–96, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017725.

Moravcsik:1989:EAR

[Mor89b] Michael J. Moravcsik. Evaluating applied research — lessons from Japan. Scientometrics, 17(1–2):181–182, July 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/BF02017732.

Motylev:1989:MPS

[Mot89] V. M. Motylev. The main problems of studying literature aging. Scientometrics, 15(1–2):97–109, January 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021802.

Moravcsik:1982:BR

[MPMG82] M. J. Moravcsik, E. Phillips, P. R. McAllister, and G. Grupp. Book reviews. Scientometrics, 4(5):399–403, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02135126.

McCain:1989:CCA

[MT89] Katherine W. McCain and Kathleen Turner. Citation context analysis and aging patterns of journal articles in molecular genetics. Scientometrics, 17(1–2):127–163, July 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017729.

Mullins:1985:ICS

[Mul85] N. C. Mullins. Invisible colleges as science elites. Scientometrics, 7(3–6):357–368, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017154.

Moed:1986:OHP

[MvR86] H. F. Moed and A. F. J. van Raan. Observations and hypotheses on the phenomenon of multiple citation to a research group’s oeuvre. Scientometrics, 10(1–2):17–33, July 1986. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016857.

Mombers:1985:DSW

[MVVL85] C. Mombers, A. Van Heeringen, R. Van Venetië, and C. Le Pair. Displaying strengths and weaknesses in national R&D performance through document cocitation. Scientometrics, 7(3– 6):341–355, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017153.

Nadel:1981:CCC

[Nad81] E. Nadel. Citation and co-citation indicators of a phased impact of the BCS theory in the physics of superconductivity. Scientometrics, 3(3):203–221, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02101666.

Nalimov:1984:LMW

[Nal84] V. V. Nalimov. Life in a multidimensional world. Scientometrics, 6(2):103–104, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02021284.

Nalimov:1987:SA

[Nal87] V. V. Nalimov. Scientists are not acrobats. Scientometrics, 12(5– 6):303–304, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02016668.

Narin:1987:BB

[Nar87] F. Narin. To believe or not to believe. Scientometrics, 12(5–6): 343–344, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02016677.

Naranan:1989:PLV

[Nar89] S. Naranan. “Power law” version of Bradford’s Law: Statistical tests and methods of estimation. Scientometrics, 17(3–4): 211–226, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026411.

Nasierowski:1988:EDM

[Nas88] W. Nasierowski. The essence and dilemmas of measurement in the sciences of organization. Scientometrics, 14(5–6):435–452, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017101.

Nederhof:1985:ERO

[Ned85] A. J. Nederhof. Evaluating research output through life work citation counts. Scientometrics, 7(1–2):23–28, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020138.

Nederhof:1988:CPP

[Ned88] A. J. Nederhof. Changes in publication patterns of biotechnologists: An evaluation of the impact of government stimulation programs in six industrial nations. Scientometrics, 14(5–6): 475–485, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017103.

Nesvetailov:1986:APS

[NG86] G. A. Nesvetailov and I. G. Gurevich. Analysis and planning of a series of multisectional scientific conferences (heat and mass transfer case). Scientometrics, 9(3–4):165–176, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017239.

Nagpaul:1989:EPC

[NG89] P. S. Nagpaul and S. P. Gupta. Effect of professional competence, managerial role and status of group leaders to R&D performance. Scientometrics, 17(3–4):301–331, September 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02026415.

Nagpaul:1988:DRP

[NK88] P. S. Nagpaul and V. S. R. Krishnaiah. Dimensions of research planning: Comparative study of research units in six countries. Scientometrics, 14(5–6):383–410, November 1988. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017098.

Nalimov:1986:BR

[NM86] V. V. Nalimov and Michael J. Moravcsik. Book reviews. Scientometrics, 9(1–2):91–97, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016610.

Narin:1985:TBS

[NN85] F. Narin and E. Noma. Is technology becoming science? Scientometrics, 7(3–6):369–381, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017155.

Noma:1982:IMA

[Nom82a] E. Noma. An improved method for analyzing square scientometric transaction matrices. Scientometrics, 4(4):297–316, July 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021645.

Noma:1982:SSC

[Nom82b] E. Noma. The simultaneous scaling of cited and citing articles in a common space. Scientometrics, 4(3):205–231, May 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021061.

Nordstrom:1987:AVB

[Nor87] L. O. Nordstrom. Applied versus basic science in the literature of plant biology: A bibliometric perspective. Scientometrics, 12(5– 6):381–393, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016681.

Nieuwenhuysen:1988:QEM

[NR88] P. Nieuwenhuysen and R. Rousseau. A quick and easy method to estimate the random effect on citation measures. Scientometrics, 13(1–2):45–52, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095762.

Nederhof:1987:CTO

[NV87a] A. J. Nederhof and A. F. J. Van Raan. Citation theory and the Ortega Hypothesis. Scientometrics, 12(5–6):325–328, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016674.

Nederhof:1987:PRB

[NV87b] A. J. Nederhof and A. F. J. Van Raan. Peer review and bibliometric indicators of scientific performance: A comparison of cum laude doctorates with ordinary doctorates in physics. Scientometrics, 11(5–6):333–350, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02279353.

Nederhof:1989:VSB

[NV89] A. J. Nederhof and A. F. J. Van Raan. A validation study of bibliometric indicators: The comparative performance of cum laude doctorates in chemistry. Scientometrics, 17(5–6):427–435, November 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017463.

Nederhof:1989:AUB

[NZDD89] A. J. Nederhof, R. A. Zwaan, R. E. De Bruin, and P. J. Dekker. Assessing the usefulness of bibliometric indicators for the humanities and the social and behavioural sciences: A comparative study. Scientometrics, 15(5–6):423–435, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017063.

Okrasa:1987:DSP

[Okr87] W. Okrasa. Differences in scientific productivity of research units: Measurement and analysis of output inequality. Scientometrics, 12(3–4):221–239, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016294.

Onodera:1988:FDF

[Ono88] N. Onodera. A frequency distribution function derived from a stochastic model considering human behaviors and its comparison with an empirical bibliometric distribution. Scientometrics, 14(1–2):143–159, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020248.

Oromaner:1981:CCR

[Oro81] M. Oromaner. Cognitive consensus in recent mainstream American sociology: An empirical analysis. Scientometrics, 3(2):73–84, March 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025631.

Oromaner:1985:OHI

[Oro85] M. Oromaner. The Ortega Hypothesis and influential articles in American sociology. Scientometrics, 7(1–2):3–10, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02020136.

Oromaner:1987:OOP

[Oro87] M. Oromaner. Ortega, obliteration and policy consequences. Scientometrics, 12(5–6):317–319, November 1987. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016672.

Oluic-Vukovic:1989:IPI

[OV89a] Vesna Oluić-Vuković. Impact of productivity increase on the distribution pattern of journals. Scientometrics, 17(1–2):97–109, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017726.

Orlov:1989:PTB

[OV89b] S. V. Orlov and A. N. Vasiljev. Possible treatment of the Bonitz– Gross effect. Scientometrics, 15(1–2):111–126, January 1989.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02021803.

Over:1988:DSI

[Ove88] R. Over. Does scholarly impact decline with age? Scientometrics, 13(5–6):215–223, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019959.

Pravdic:1988:SNC

[PABK88] Nevenka Pravdić, Alena Aganović-Boras, and Dubravka Kritovac. In search of a “non-citation index” indicator for scientific activity assessment in less developed countries. Case study of Croatia/ Yugoslavia. Scientometrics, 14(1–2):111–125, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020246.

Pavitt:1985:PSI

[Pav85] K. Pavitt. Patent statistics as indicators of innovative activities: Possibilities and problems. Scientometrics, 7(1–2):77–99, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020142.

Pavitt:1988:SSB

[Pav88] K. Pavitt. The size and structure of British technology activities: What we do and do not know. Scientometrics, 14(3–4): 329–346, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020083.

Porter:1985:ICD

[PC85] A. L. Porter and D. E. Chubin. An indicator of cross-disciplinary research. Scientometrics, 8(3–4):161–176, September 1985. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016934.

Porter:1988:CSP

[PCJ88] A. L. Porter, D. E. Chubin, and Xiao-Yin Jin. Citations and scientific progress: Comparing bibliometric measures with scientist judgments. Scientometrics, 13(3–4):103–124, March 1988.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017178.

Peritz:1983:MPM

[Per83a] B. C. Peritz. Are methodological papers more cited than theoretical or empirical ones? The case of sociology. Scientometrics, 5(4):211–218, July 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019738.

Peritz:1983:CCR

[Per83b] Bluma C. Peritz. A classification of citation roles for the social sciences and related fields. Scientometrics, 5(5):303–312, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02147226.

Persson:1986:OBR

[Per86] O. Persson. Online bibliometrics. A research tool for every man. Scientometrics, 10(1–2):69–75, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016861.

Pinski:1980:CBM

[Pin80] G. Pinski. Citation based measures of research interactivity. Scientometrics, 2(4):257–263, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016347.

Puzikov:1987:QEB

[PK87] M. D. Puzikov and A. E. Kasjanov. Quantitative estimation of “big” and “little” science interrelation. Scientometrics, 11(1– 2):99–104, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016633.

Pontigo:1986:QAB

[PL86] J. Pontigo and F. W. Lancaster. Qualitative aspects of the Bradford distribution. Scientometrics, 9(1–2):59–70, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016608.

Prabha:1987:CSC

[PL87] C. G. Prabha and F. W. Lancaster. Comparing the scatter of citing and cited literature. Scientometrics, 12(1–2):17–31, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016687.

Plomp:1989:SRC

[Plo89] R. Plomp. Statistical reliability of citation frequency as an indicator of scientific impact. Scientometrics, 17(1–2):71–81, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017724.

Pruthi:1987:IRP

[PNN87] S. Pruthi, P. S. Nagpaul, and S. A. Nabi. Indicators of research planning: A comparative study of research groups in six countries. Scientometrics, 11(3–4):141–161, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016588.

Poikolainen:1984:OFM

[Poi84] Kari Poikolainen. Organization and funding of medical research in 10 European countries. Scientometrics, 6(5):327–358, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02020131.

Pokrovsky:1980:SPM

[Pok80] V. A. Pokrovsky. Some problems of measuring the impact of R&D upon the efficiency of social production. Scientometrics, 2 (2):121–132, March 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02279363.

Porteous:1988:RDQ

[Por88] M. Porteous. The role and development of quantitative indicators for research and technology policy making: Some experience from the Department of Trade and Industry. Scientometrics, 14(3–4): 315–327, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020082.

Pouris:1989:SAA

[Pou89] A. Pouris. A scientometric assessment of agricultural research in South Africa. Scientometrics, 17(5–6):401–413, November 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017461.

Pravdic:1986:DAM

[POV86] Nevenka Pravdić and Vesna Oluić-Vuković. Dual approach to multiple authorship in the study of collaboration/scientific output relationship. Scientometrics, 10(5–6):259–280, November 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016774.

Pravdic:1985:CPA

[PP85] Nevenka Pravdić and R. Pekorari. The citing practices of the authors to the national journals in mathematics, physics, and chemistry. Scientometrics, 8(3–4):233–246, September 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016938.

Pyenson:1984:PPW

[PS84] L. Pyenson and M. Singh. Physics on the periphery: A world survey, 1920–1929. Scientometrics, 6(5):279–306, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02020129.

Phillips:1988:BUS

[PT88] D. C. Phillips and J. Turney. Bibliometrics and UK science policy. Scientometrics, 14(3–4):185–200, September 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020074.

Purica:1988:CIS

[Pur88] I. I. Purica. Creativity, intelligence and synergetic processes in the development of science. Scientometrics, 13(1–2):11–24, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095759.

Purica:1989:CSC

[Pur89] I. I. Purica. Creativity and the socio-cultural niche. Scientometrics, 15(3–4):181–187, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017197.

Qurashi:1984:PRF

[Qur84] M. M. Qurashi. Publication rate as a function of the laboratory/ group size. Scientometrics, 6(1):19–26, January 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020110.

Rajeswari:1983:QAI

[Raj83] Mrs A. R. Rajeswari. A quantitative analysis of Indian science and technology manpower employment and economic development. Scientometrics, 5(6):343–359, November 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02134323.

Rajeswari:1989:FST

[Raj89] A. R. Rajeswari. Forecasting of science & technology expenditure of India by simulation method. Scientometrics, 17(3–4): 227–251, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026412.

Rice:1989:JJC

[RBBH89] R. E. Rice, Christine L. Borgman, Diane Bednarski, and P. J. Hart. Journal-to-journal citation data: Issues of validity and reliability. Scientometrics, 15(3–4):257–282, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017202.

Rip:1984:CWM

[RC84] A. Rip and J.-P. Courtial. Co-word maps of biotechnology: An example of cognitive scientometrics. Scientometrics, 6(6): 381–400, November 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02025827.

Rabkin:1983:BR

[RFH^{+}83] Yakov M. Rabkin, J. Davidson Frame, S. D. Haitun, A. V. Pilipenko, and Michael J. Moravcsik. Book reviews. Scientometrics, 5(3):195–203, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095629.

Richards:1984:SSA

[Ric84] J. M. Richards, Jr. Structure of specialization among American population scientists. Scientometrics, 6(6):425–432, November 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025829.

Riecken:1980:VSB

[Rie80] H. W. Riecken. Vital signs for basic research in the behavioral and social sciences. Scientometrics, 2(5–6):435–437, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095087.

Rothman:1985:UBI

[RL85] H. Rothman and G. Lester. The use of bibliometric indicators in the study of insecticide research. Scientometrics, 8(3–4): 247–262, September 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016939.

Rushton:1981:RPU

[RM81] J. P. Rushton and S. Meltzer. Research productivity, university revenue, and scholarly impact (citations) of 169 British, Canadian and United States universities (1977). Scientometrics, 3(4):275– 303, July 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021122.

Riley:1986:BR

[RMD86] Kathryn Riley, Robert P. Multhauf, and A. Donovan. Book reviews. Scientometrics, 10(3–4):221–225, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02026043.

Russell:1987:PLC

[RMM87] Jane M. Russell, M. Mendoza, and G. Martinez. Patterns of literature citation by undergraduate students and researchers in the veterinary field. Scientometrics, 12(1–2):73–80, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016690.

Rushton:1983:PRC

[RMP83] J. P. Rushton, H. G. Murray, and S. V. Paunonen. Personality, research creativity, and teaching effectiveness in university professors. Scientometrics, 5(2):93–116, March 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02072856.

Roche:1983:SC

[Roc83] Marcel Roche. Sciences and cultures. Scientometrics, 5(5): 329–331, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02147228.

Rosenberg:1980:CII

[Ros80] Nathan Rosenberg. Comments on “Indicators of the impact of R&D on the economy”. Scientometrics, 2(5–6):387–393, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095081. See [Fre80].

Rousseau:1987:GTU

[Rou87] R. Rousseau. The Gozinto Theorem: Using citations to determine influences on a scientific publication. Scientometrics, 11(3– 4):217–229, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016593.

Rousseau:1989:MDS

[Rou89] R. Rousseau. Merging data sets. Scientometrics, 15(3–4):305–308, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017205.

Roy:1983:ATC

[RRJ83] Rustum Roy, N. R. Roy, and G. G. Johnson, Jr. Approximating total citation counts from first author counts and from total papers. Scientometrics, 5(2):117–124, March 1983. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02072857.

Schubert:1981:SSM

[SB81] A. Schubert and T. Braun. Some scientometric measures of publishing performance for 85 Hungarian research institutes. Scientometrics, 3(5):379–388, September 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017575.

Schubert:1986:RIR

[SB86a] A. Schubert and T. Braun. Relative indicators and relational charts for comparative assessment of publication output and citation impact. Scientometrics, 9(5–6):281–291, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017249.

Smart:1986:ACI

[SB86b] J. C. Smart and A. E. Bayer. Author collaboration and impact: A note on citation rates of single and multiple authored articles. Scientometrics, 10(5–6):297–305, November 1986. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016776.

Studer:1984:SAU

[SBN84] K. E. Studer, E. J. Barboni, and K. B. Numan. Structural analysis using the input-output model: With special reference to networks of science. Scientometrics, 6(6):401–423, November 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02025828.

Schubert:1983:QSSa

[Sch83a] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 5(2):125–133, March 1983. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02072858.

Schubert:1983:QSSb

[Sch83b] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 5(3):189–194, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095628.

Schubert:1983:QSSc

[Sch83c] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 5(6):397–403, November 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02134326.

Schubert:1984:QSSa

[Sch84a] A. Schubert. Quantitative studies of science. A current bibliography. Scientometrics, 6(1):55–59, January 1984. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020113.

Schubert:1984:QSSb

[Sch84b] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 6(3):197–202, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016763.

Schubert:1984:QSSc

[Sch84c] A. Schubert. Quantitative studies of science. A current bibliography. Scientometrics, 6(5):359–367, September 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020132.

Schubert:1985:QSSa

[Sch85a] A. Schubert. Quantitative studies of science. A current bibliography. Scientometrics, 7(1–2):111–117, January 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020145.

Schubert:1985:QSSb

[Sch85b] A. Schubert. Quantitative studies of science: a current bibliography. Scientometrics, 8(1–2):137–140, July 1985. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02025226.

Schubert:1986:QSS

[Sch86] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 9(5–6):293–304, May 1986. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017250.

Schubert:1987:QSS

[Sch87] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 12(5–6):395–412, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016682.

Schubert:1988:QSSa

[Sch88a] A. Schubert. Quantitative studies of science: a current bibliography. Scientometrics, 13(3–4):139–172, March 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017181.

Schubert:1988:QSSb

[Sch88b] A. Schubert. Quantitative studies of science: a current bibliography. Scientometrics, 14(5–6):513–520, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017106.

Schubert:1989:QSSa

[Sch89a] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 17(1–2):171–180, July 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017731.

Schubert:1989:QSSb

[Sch89b] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 17(3–4):373–380, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02026419.

Snizek:1982:SPP

[SDH82] W. E. Snizek, C. J. Dudley, and J. E. Hughes. The second process of peer review: Some correlates of comments published in the ASR (1947–1979). Scientometrics, 4(6):417–430, November 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021138.

Smart:1982:CFS

[SE82] J. C. Smart and C. F. Elton. Consumption factor scores of psychology journals: Scientometric properties and qualitative implications. Scientometrics, 4(5):349–360, September 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02135121.

Sengupta:1984:PPS

[Sen84] I. N. Sengupta. The place of phenomenological studies in scientometrics. Scientometrics, 6(2):109–113, March 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021286.

Sengupta:1985:GBL

[Sen85] I. N. Sengupta. The growth of biophysical literature. Scientometrics, 8(5–6):365–375, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018059.

Sengupta:1986:TNP

[Sen86a] I. N. Sengupta. Three new parameters in bibliometric research and their application to rerank periodicals in the field of biochemistry. Scientometrics, 10(5–6):235–242, November 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016772.

Senter:1986:CMP

[Sen86b] R. Senter, Jr. A causal model of productivity in a research facility. Scientometrics, 10(5–6):307–328, November 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016777.

Sen:1989:BSG

[Sen89a] S. K. Sen. Bibliographic scattering: A generalised source approach. Scientometrics, 17(3–4):197–204, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02026409.

Sen:1989:NTC

[Sen89b] S. K. Sen. A note on theoretical correlation between Bradford’s Law and recently proposed linear equation of the type R(r) = a.r −b. Scientometrics, 17(3–4):205–210, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02026410.

Sengupta:1989:GKL

[Sen89c] I. N. Sengupta. The growth of knowledge and literature in neuroscience. Scientometrics, 17(3–4):253–288, September 1989.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02026413.

Sengupta:1989:WFR

[Sen89d] I. N. Sengupta. A weightage formula to rerank periodicals in the field of microbiology. Scientometrics, 17(3–4):289–300, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02026414.

Self:1989:AIS

[SFL89] Ph. C. Self, Th. W. Filardo, and F. W. Lancaster. Acquired immunodeficiency syndrome (AIDS) and the epidemic growth of its literature. Scientometrics, 17(1–2):49–60, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017722.

Small:1980:CCA

[SG80] H. Small and E. Greenlee. Citation context analysis of a cocitation cluster: Recombinant–DNA. Scientometrics, 2(4):277– 301, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016349.

Schubert:1983:SRC

[SG83] A. Schubert and W. Glänzel. Statistical reliability of comparisons based on the citation impact of scientific publications. Scientometrics, 5(1):59–73, January 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02097178.

Schubert:1984:DLC

[SG84] A. Schubert and W. Glänzel. A dynamic look at a class of skew distributions. A model with scientometric applications. Scientometrics, 6(3):149–167, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016759.

Small:1986:CR

[SG86] H. Small and E. Greenlee. Collagen research in the 1970s. Scientometrics, 10(1–2):95–117, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016863.

Schubert:1987:SFC

[SGB87] A. Schubert, W. Glänzel, and T. Braun. Subject field characteristic citation scores and scales for assessing research performance. Scientometrics, 12(5–6):267–291, November 1987. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016664.

Schubert:1989:SDC

[SGB89] A. Schubert, W. Glänzel, and T. Braun. Scientometric datafiles. A comprehensive set of indicators on 2649 journals and 96 countries in all major science fields and subfields 1981–1985. Scientometrics, 16(1–6):3–478, June 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02093234.

Snizek:1983:EAV

[SH83] W. E. Snizek and M. Hughes. An empirical assessment of the validity of Mullins’ theory group classifications. Scientometrics, 5(3):155–162, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095625.

Shaw:1981:ITS

[Sha81] W. M. Shaw, Jr. Information theory and scientific communication. Scientometrics, 3(3):235–249, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02101668.

Shaw:1987:AAC

[Sha87] J. G. Shaw. Article-by-article citation analysis of medical journals. Scientometrics, 12(1–2):101–110, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016692.

Sharabchiev:1989:CAB

[Sha89] J. T. Sharabchiev. Cluster analysis of bibliographic references as a scientometric method. Scientometrics, 15(1–2):127–137, January 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02021804.

Shenhav:1989:CAT

[SHC89] Y. A. Shenhav, Y. Haberfeld, and B. P. Cohen. Contextual analysis of team productivity in the R&D industry. Scientometrics, 17(5–6):387–400, November 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017460.

Sheldon:1980:CTP

[She80] J. C. Sheldon. A cybernetic theory of physical science professions: The causes of periodic normal and revolutionary science between 1000 and 1870AD. Scientometrics, 2(2):147–167, March 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02279365.

Shelishch:1982:QSB

[She82] P. B. Shelishch. A quantitative study of biologists in the 18th and 19th centuries. Scientometrics, 4(4):317–329, July 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021646.

Shrum:1985:QJT

[Shr85] W. Shrum. Quality judgments of technical fields: Bias, marginality, and the role of the elite. Scientometrics, 8(1–2):35–57, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025220.

Simonton:1980:TSA

[Sim80] D. K. Simonton. Techno-scientific activity and war: A yearly time-series analysis, 1500–1903 A.D. Scientometrics, 2(4):251– 255, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016346.

Simonton:1984:SEH

[Sim84] D. K. Simonton. Scientific eminence historical and contemporary: A measurement assessment. Scientometrics, 6(3):169–182, May 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016760.

Simonton:1986:MDS

[Sim86a] D. K. Simonton. Multiple discovery: Some Monte Carlo simulations and Gedanken experiments. Scientometrics, 9(5–6):269–280, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017248.

Simonton:1986:MPD

[Sim86b] D. K. Simonton. Multiples, Poisson distributions, and chance: An analysis of the Brannigan–Wanner model. Scientometrics, 9 (3–4):127–137, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017236.

Sitarska:1987:SBW

[Sit87] Anna Sitarska. Scientometrics and bibliometrics in the Warsaw University curriculum of library and information science: Place and field structure. Scientometrics, 12(3–4):241–257, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016295.

Sen:1986:BEL

[SK86] Subir K. Sen and Rameshi Kundra. Bibliometrics of English language alcohol fuel literature. A new empirical equation of scatter. Scientometrics, 10(1–2):43–54, July 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016859.

Singh:1989:AWC

[SK89] Prithpal Singh and V. S. R. Krishnaiah. Analysis of work climate perceptions and performance of research and development units. Scientometrics, 17(3–4):333–351, September 1989. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02026416.

Sullivan:1980:URT

[SKWK80] D. Sullivan, D. Koester, D. H. White, and R. Kern. Understanding rapid theoretical change in particle physics: A month-by-month co-citation analysis. Scientometrics, 2(4):309–319, July 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016351.

Slater:1983:HCM

[Sla83] P. B. Slater. Hierarchical clustering of mathematical journals based upon citation matrices. Scientometrics, 5(1):55–58, January 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02097177.

Schubert:1983:BR

[SM83] András Schubert and Michael J. Moravcsik. Book reviews. Scientometrics, 5(2):135–137, March 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02072859.

Small:1987:SBR

[Sma87] H. Small. The significance of bibliographic references. Scientometrics, 12(5–6):339–341, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016676.

Small:1988:MDS

[Sma88] H. Small. Mapping the dynamics of science and technology. Scientometrics, 14(1–2):165–168, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020250.

Simeon:1986:MSA

[SMKG86] Vl. Simeon, B. Momčilović, Z. Kralj, and B. Grgas. Multivariate statistical analysis of the bibliographic output from a research institution, in relation to the measures of scientific policy. Scientometrics, 9(5–6):223–230, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017245.

Schubert:1984:BR

[SMR84] A. Schubert, Michael J. Moravcsik, and Sal Restivo. Book reviews. Scientometrics, 6(6):445–448, November 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02025831.

Snizek:1984:CFR

[Sni84] W. E. Snizek. Casting the first rock: Some observations on the bestowers and recipients of journal article comments in sociology. Scientometrics, 6(4):215–222, July 1984. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02279356.

Snizek:1986:REO

[Sni86] W. E. Snizek. A re-examination of the Ortega Hypothesis: The Dutch case. Scientometrics, 9(1–2):3–11, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02016603.

Snizek:1987:SIT

[Sni87] W. E. Snizek. In search of influence: The testing of the Ortega Hypothesis. Scientometrics, 12(5–6):311–314, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016670.

Satyanarayana:1989:APL

[SR89] K. Satyanarayana and K. V. Ratnakar. Authorship patterns in life sciences, preclinical basic and clinical research papers. Scientometrics, 17(3–4):363–371, September 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02026418.

Small:1985:CSCa

[SS85] H. Small and E. Sweeney. Clustering the Science Citation Index^{(}R) using co-citations. Scientometrics, 7(3–6):391–409, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017157.

Small:1985:CSCb

[SSG85] H. Small, E. Sweeney, and E. Greenlee. Clustering the science citation index using co-citations. II. Mapping science. Scientometrics, 8(5–6):321–340, November 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018057.

Schubert:1986:PPI

[ST86] A. Schubert and A. Telcs. Publication potential — an indicator of scientific strength for cross-national comparisons. Scientometrics, 9(5–6):231–238, May 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017246.

Stern:1981:N

[Ste81] L. Stern. News. Scientometrics, 3(3):253–259, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02101670.

Stefaniak:1982:IMA

[Ste82] B. Stefaniak. Individual and multiple authorship of papers in chemistry and physics. Scientometrics, 4(4):331–337, July 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Stefaniak:1985:PLI

[Ste85] Barbara Stefaniak. Periodical literature of information science as reflected in Referativnyj Zhurnal, section 59, informatika. Scientometrics, 7(3–6):177–194, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017145.

Stefaniak:1987:UBD

[Ste87] Barbara Stefaniak. Use of bibliographic data bases for scientometric studies. Scientometrics, 12(3–4):149–161, September 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02016289.

Soete:1983:UFP

[SW83] L. G. Soete and Sally M. E. Wyatt. The use of foreign patenting as an internationally comparable science and technology output indicator. Scientometrics, 5(1):31–54, January 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02097176.

Swales:1985:ELP

[Swa85] J. M. Swales. English language papers and authors’ first language: Preliminary explorations. Scientometrics, 8(1–2):91–101, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02025223.

Swales:1988:LSC

[Swa88] J. Swales. Language and scientific communication: The case of the reprint request. Scientometrics, 13(3–4):93–101, March 1988.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017177.

Szabo:1985:ACE

[Sza85] A. T. Szabó. Alphonse de Candolle’s early scientometrics (1883, 1885) with references to recent trends in the field (1978–1983). Scientometrics, 8(1–2):13–33, July 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02025219.

Schubert:1983:SAA

[SZB83] A. Schubert, S. Zsindely, and T. Braun. Scientometric analysis of attendance at international scientific meetings. Scientometrics, 5(3):177–187, May 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095627.

Schubert:1985:SIE

[SZB85] A. Schubert, S. Zsindely, and T. Braun. Scientometric indicators for evaluating medical research output of mid-size countries. Scientometrics, 7(3–6):155–163, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017143.

Schubert:1984:QAV

[SZTB84] A. Schubert, S. Zsindely, A. Telcs, and T. Braun. Quantitative analysis of a visible tip of the peer review iceberg: Book reviews in chemistry. Scientometrics, 6(6):433–443, November 1984. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02025830.

Todorov:1986:FCS

[TA86] R. Todorov and K. Atanassov. Formal communication in science: A model based on generalized nets. Scientometrics, 9(3– 4):177–185, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017240.

Tijseen:1987:QCA

[TDV87] R. J. W. Tijseen, J. De Leeuw, and A. F. J. Van Raan. Quasicorrespondence analysis on scientometric transaction matrices.

Scientometrics, 11(5–6):351–366, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02279354.

Turner:1981:IBR

[TK81] C. F. Turner and Sara B. Kiesler. The impact of basic research in the social sciences: The case of education. Scientometrics, 3(3):177–190, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02101664.

Todorov:1983:CMP

[Tod83] R. Todorov. Condensed matter physics journals. Scientometrics, 5(5):291–301, September 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02147225.

Todorov:1984:DIW

[Tod84] R. Todorov. Determination of influence weights for scientific journals. Scientometrics, 6(2):127–137, March 1984. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021288.

Todorov:1985:DPL

[Tod85] R. Todorov. Distribution of physics literature. Scientometrics, 7 (3–6):195–209, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017146.

Todorov:1989:RSF

[Tod89] R. Todorov. Representing a scientific field: A bibliometric approach. Scientometrics, 15(5–6):593–605, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017072.

Trofimenko:1987:SAD

[Tro87] A. P. Trofimenko. Scientometric analysis of the development of nuclear physics during the last 50 years. Scientometrics, 11(3– 4):231–250, March 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016594.

Tijssen:1989:MCW

[TV89] R. J. W. Tijssen and A. F. J. Van Raan. Mapping co-word structures: A comparison of multidimensional scaling and leximappe. Scientometrics, 15(3–4):283–295, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02017203.

VanHeeringen:1981:DRG

[Van81] A. Van Heeringen. Dutch research groups: Output and collaboration. Scientometrics, 3(4):305–315, July 1981. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02021123.

VanRaan:1988:CHS

[Van88] A. F. J. Van Raan. Comments on Henry Small, recipient of the 1987 Derek de Solla Price Award. Scientometrics, 14(5–6): 361–363, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017096. See [Ano88b].

VanRaan:1989:CFN

[Van89a] A. F. J. Van Raan. Comments on Francis Narin, recipient of the 1988 Derek de Solla Price Award. Scientometrics, 17(1–2):5–7, July 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017717.

VanRossum:1989:ODP

[Van89b] W. Van Rossum. Operationalizing developments in a problem field: The case of MBD. Scientometrics, 15(5–6):509–526, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02017068.

VanHeeringen:1987:RBAa

[VD87a] A. Van Heeringen and P. A. Dijkwel. The relationships between age, mobility and scientific productivity. Part I. Scientometrics, 11(5–6):267–280, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02279349.

VanHeeringen:1987:RBAb

[VD87b] A. Van Heeringen and P. A. Dijkwel. The relationships between age, mobility and scientific productivity. Part II. Scientometrics, 11(5–6):281–293, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02279350.

Velho:1986:MCC

[Vel86] Léa Velho. The “meaning” of citation in the context of a scientifically peripheral country. Scientometrics, 9(1–2):71–89, January 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016609.

Velho:1987:ABH

[Vel87] Lea Velho. The author and the beholder: How paradigm commitments can influence the interpretation of research results. Scientometrics, 11(1–2):59–70, January 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02016630.

vanHouten:1983:MPO

[vHvVLD83] J. van Houten, H. G. van Vuren, C. Le Pair, and G. Dijkhuis. Migration of physicists to other academic disciplines: Situation in the Netherlands. Scientometrics, 5(4):257–267, July 1983. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02019741.

Vinkler:1986:ESM

[Vin86] P. Vinkler. Evaluation of some methods for the relative assessment of scientific publications. Scientometrics, 10(3–4):157– 177, September 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026039.

Vinkler:1987:QQC

[Vin87] P. Vinkler. A quasi-quantitative citation model. Scientometrics, 12(1–2):47–72, July 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016689.

Vinkler:1988:ASC

[Vin88a] P. Vinkler. An attempt of surveying and classifying bibliometric indicators for scientometric purposes. Scientometrics, 13(5– 6):239–259, May 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019961.

Vinkler:1988:BFS

[Vin88b] P. Vinkler. Bibliometric features of some scientific subfields and the scientometric consequences therefrom. Scientometrics, 14(5– 6):453–474, November 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017102.

Vinkler:1988:WIP

[Vin88c] P. Vinkler. Weighted impact of publications and relative contribution score. Two new indicators characterizing publication activity of countries. Scientometrics, 14(1–2):161–163, July 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02020249.

VanEls:1989:CGB

[VJL89] W. P. Van Els, C. N. M. Jansz, and C. Le Pair. The citation gap between printed and instrumental output of technological research: The case of the electron microscope. Scientometrics, 17(5– 6):415–425, November 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017462.

Vlachy:1985:CHS

[Vla85] J. Vlachý. Citation histories of scientific publications. The data sources. Scientometrics, 7(3–6):505–528, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017164.

VanRaan:1989:DSF

[VP89] A. F. J. Van Raan and H. P. F. Peters. Dynamics of a scientific field analysed by co-subfield structures. Scientometrics, 15 (5–6):607–620, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017073.

VanSteijn:1988:RTL

[VR88] F. Van Steijn and A. Rip. The role of trade literature in the communication system. Scientometrics, 13(3–4):81–91, March 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017176.

vanRaan:1987:CIS

[vRH87] A. F. J. van Raan and D. Hartmann. The comparative impact of scientific publications and journals: Methods of measurement and graphical display. Scientometrics, 11(5–6):325–331, May 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02279352.

Wood:1988:GSO

[Woo88] Judith B. Wood. The growth of scholarship: An online bibliometric comparison of dissertations in the sciences and humanities. Scientometrics, 13(1–2):53–62, January 1988. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095763.

Yablonsky:1980:FRD

[Yab80] A. I. Yablonsky. On fundamental regularities of the distribution of scientific productivity. Scientometrics, 2(1):3–34, January 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016597.

Yablonsky:1985:SNG

[Yab85] A. I. Yablonsky. Stable non-Gaussian distributions in scientometrics. Scientometrics, 7(3–6):459–470, March 1985. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/BF02017161.

Yamazaki:1989:RSE

[Yam89] S. Yamazaki. Referee systems of English-language scientific journals in Japan. Scientometrics, 15(3–4):297–303, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017204.

Yanovsky:1981:CAS

[Yan81] V. I. Yanovsky. Citation analysis significance of scientific journals. Scientometrics, 3(3):223–233, May 1981. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02101667.

Yankevich:1982:API

[Yan82] W. F. Yankevich. Analysis of publication and invention productivity in some Soviet academic institutions. Scientometrics, 4(6): 431–437, November 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02021139.

Yaalon:1989:BR

[YT89] Dan H. Yaalon and Jean Tague. Book reviews. Scientometrics, 15 (3–4):313–315, March 1989. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017207.

Yuthavong:1986:BIS

[Yut86] Yongyuth Yuthavong. Bibliometric indicators of scientific activity in Thailand. Scientometrics, 9(3–4):139–143, March 1986. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02017237.

Zuckerman:1980:ISN

[ZM80a] Harriet Zuckerman and Roberta Balstad Miller. Indicators of science: Notes and queries. Scientometrics, 2(5–6):347–353, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02095077.

Zuckerman:1980:I

[ZM80b] Harriet Zuckerman and Roberta Balstad Miller. Introduction. Scientometrics, 2(5–6):327–330, October 1980. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02095074.

Zsindely:1982:CPE

[ZSB82a] S. Zsindely, A. Schubert, and T. Braun. Citation patterns of editorial gatekeepers in international chemistry journals. Scientometrics, 4(1):69–76, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02098007.

Zsindely:1982:EGP

[ZSB82b] S. Zsindely, A. Schubert, and T. Braun. Editorial gatekeeping patterns in international science journals. A new science indicator. Scientometrics, 4(1):57–68, January 1982. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1007/BF02098006.

Zuckerman:1987:CAC

[Zuc87] Harriet Zuckerman. Citation analysis and the complex problem of intellectual influence. Scientometrics, 12(5–6):329–338, November 1987. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/article/
10.1007/BF02016675.